### v3

# Matryoshka Representation Learning

Abstract
Learned representations are a central component in modern ML systems, serving a multitude of downstream tasks. When training such representations, it is often the case that computational and statistical constraints for each downstream task are unknown. In this context, rigid fixed-capacity representations can be either over or under-accommodating to the task at hand. This leads us to ask: can we design a flexible representation that can adapt to multiple downstream tasks with varying computational resources? Our main contribution is [Uncaptioned image] $\rm Matryoshka~Representation~Learning$} ($\rm MRL$}) which encodes information at different granularities and allows a single embedding to adapt to the computational constraints of downstream tasks. $\rm MRL$} minimally modifies existing representation learning pipelines and imposes no additional cost during inference and deployment. $\rm MRL$} learns coarse-to-fine representations that are at least as accurate and rich as independently trained low-dimensional representations. The flexibility within the learned $\rm Matryoshka~Representations$} offer: (a) up to $\mathbf{14}\times$ smaller embedding size for ImageNet-1K classification at the same level of accuracy; (b) up to $\mathbf{14}\times$ real-world speed-ups for large-scale retrieval on ImageNet-1K and 4K; and (c) up to $\mathbf{2}$% accuracy improvements for long-tail few-shot classification, all while being as robust as the original representations. Finally, we show that $\rm MRL$} extends seamlessly to web-scale datasets (ImageNet, JFT) across various modalities – vision (ViT, ResNet), vision + language (ALIGN) and language (BERT). $\rm MRL$} code and pretrained models are open-sourced at https://github.com/RAIVNLab/MRL.
1 Introduction
Learned representations [55] are fundamental building blocks of real-world ML systems [62, 86]. Trained once and frozen, $d$-dimensional representations encode rich information and can be used to perform multiple downstream tasks [4]. The deployment of deep representations has two steps: (1) an expensive yet constant-cost forward pass to compute the representation [27] and (2) utilization of the representation for downstream applications [48, 84]. Compute costs for the latter part of the pipeline scale with the embedding dimensionality as well as the data size ($N$) and label space ($L$). At web-scale [15, 80] this utilization cost overshadows the feature computation cost. The rigidity in these representations forces the use of high-dimensional embedding vectors across multiple tasks despite the varying resource and accuracy constraints that require flexibility.
Human perception of the natural world has a naturally coarse-to-fine granularity [26, 30]. However, perhaps due to the inductive bias of gradient-based training [79], deep learning models tend to diffuse “information” across the entire representation vector. The desired elasticity is usually enabled in the existing flat and fixed representations either through training multiple low-dimensional models [27], jointly optimizing sub-networks of varying capacity [9, 95] or post-hoc compression [36, 57]. Each of these techniques struggle to meet the requirements for adaptive large-scale deployment either due to training/maintenance overhead, numerous expensive forward passes through all of the data, storage and memory cost for multiple copies of encoded data, expensive on-the-fly feature selection or a significant drop in accuracy. By encoding coarse-to-fine-grained representations, which are as accurate as the independently trained counterparts, we learn with minimal overhead a representation that can be deployed adaptively at no additional cost during inference.
We introduce [Uncaptioned image] $\rm Matryoshka~Representation~Learning$} ($\rm MRL$}) to induce flexibility in the learned representation. $\rm MRL$} learns representations of varying capacities within the same high-dimensional vector through explicit optimization of $O(\log(d))$ lower-dimensional vectors in a nested fashion, hence the name $\rm Matryoshka$}. $\rm MRL$} can be adapted to any existing representation pipeline and is easily extended to many standard tasks in computer vision and natural language processing. Figure 1 illustrates the core idea of $\rm Matryoshka~Representation~Learning$} ($\rm MRL$}) and the adaptive deployment settings of the learned $\rm Matryoshka~Representations$}.
Figure 1: Refer to caption $\rm Matryoshka~Representation~Learning$} is adaptable to any representation learning setup and begets a $\rm Matryoshka~Representation$} $z$ by optimizing the original loss $\mathcal{L}(.)$ at $O(\log(d))$ chosen representation sizes. $\rm Matryoshka~Representation$} can be utilized effectively for adaptive deployment across environments and downstream tasks.
The first $m$-dimensions, $m\in[d]$, of the $\rm Matryoshka~Representation$} is an information-rich low-dimensional vector, at no additional training cost, that is as accurate as an independently trained $m$-dimensional representation. The information within the $\rm Matryoshka~Representation$} increases with the dimensionality creating a coarse-to-fine grained representation, all without significant training or additional deployment overhead. $\rm MRL$} equips the representation vector with the desired flexibility and multifidelity that can ensure a near-optimal accuracy-vs-compute trade-off. With these advantages, $\rm MRL$} enables adaptive deployment based on accuracy and compute constraints.
The $\rm Matryoshka~Representations$} improve efficiency for large-scale classification and retrieval without any significant loss of accuracy. While there are potentially several applications of coarse-to-fine $\rm Matryoshka~Representations$}, in this work we focus on two key building blocks of real-world ML systems: large-scale classification and retrieval. For classification, we use adaptive cascades with the variable-size representations from a model trained with $\rm MRL$}, significantly reducing the average dimension of embeddings needed to achieve a particular accuracy. For example, on ImageNet-1K, $\rm MRL$} + adaptive classification results in up to a $14\times$ smaller representation size at the same accuracy as baselines (Section 4.2.1). Similarly, we use $\rm MRL$} in an adaptive retrieval system. Given a query, we shortlist retrieval candidates using the first few dimensions of the query embedding, and then successively use more dimensions to re-rank the retrieved set. A simple implementation of this approach leads to $128\times$ theoretical (in terms of FLOPS) and $14\times$ wall-clock time speedups compared to a single-shot retrieval system that uses a standard embedding vector; note that $\rm MRL$}’s retrieval accuracy is comparable to that of single-shot retrieval (Section 4.3.1). Finally, as $\rm MRL$} explicitly learns coarse-to-fine representation vectors, intuitively it should share more semantic information among its various dimensions (Figure 5). This is reflected in up to $2\%$ accuracy gains in long-tail continual learning settings while being as robust as the original embeddings. Furthermore, due to its coarse-to-fine grained nature, $\rm MRL$} can also be used as method to analyze hardness of classification among instances and information bottlenecks.
We make the following key contributions:
  1. 1.
    We introduce [Uncaptioned image] $\rm Matryoshka~Representation~Learning$} ($\rm MRL$}) to obtain flexible representations ($\rm Matryoshka~Representations$}) for adaptive deployment (Section 3).
  2. 2.
    Up to $14\times$ faster yet accurate large-scale classification and retrieval using $\rm MRL$} (Section 4).
  3. 3.
    Seamless adaptation of $\rm MRL$} across modalities (vision - ResNet & ViT, vision + language - ALIGN, language - BERT) and to web-scale data (ImageNet-1K/4K, JFT-300M and ALIGN data).
  4. 4.
    Further analysis of $\rm MRL$}’s representations in the context of other downstream tasks (Section 5).
2 Related Work
Representation Learning.
Large-scale datasets like ImageNet [16, 71] and JFT [80] enabled the learning of general purpose representations for computer vision [4, 93]. These representations are typically learned through supervised and un/self-supervised learning paradigms. Supervised pretraining [27, 49, 77] casts representation learning as a multi-class/label classification problem, while un/self-supervised learning learns representation via proxy tasks like instance classification [92] and reconstruction [29, 60]. Recent advances [12, 28] in contrastive learning [25] enabled learning from web-scale data [21] that powers large-capacity cross-modal models [18, 44, 67, 96]. Similarly, natural language applications are built [38] on large language models [8] that are pretrained [64, 70] in a un/self-supervised fashion with masked language modelling [19] or autoregressive training [66].
[Uncaptioned image] $\rm Matryoshka~Representation~Learning$} ($\rm MRL$}) is complementary to all these setups and can be adapted with minimal overhead (Section 3). $\rm MRL$} equips representations with multifidelity at no additional cost which enables adaptive deployment based on the data and task (Section 4).
Efficient Classification and Retrieval.
Efficiency in classification and retrieval during inference can be studied with respect to the high yet constant deep featurization costs or the search cost which scales with the size of the label space and data. Efficient neural networks address the first issue through a variety of algorithms [24, 52] and design choices [37, 51, 82]. However, with a strong featurizer, most of the issues with scale are due to the linear dependence on number of labels ($L$), size of the data ($N$) and representation size ($d$), stressing RAM, disk and processor all at the same time.
The sub-linear complexity dependence on number of labels has been well studied in context of compute [3, 41, 65] and memory [20] using Approximate Nearest Neighbor Search (ANNS) [59] or leveraging the underlying hierarchy [17, 53]. In case of the representation size, often dimensionality reduction [72, 83], hashing techniques [14, 50, 73] and feature selection [61] help in alleviating selective aspects of the $O(d)$ scaling at a cost of significant drops in accuracy. Lastly, most real-world search systems [11, 15] are often powered by large-scale embedding based retrieval [10, 62] that scales in cost with the ever increasing web-data. While categorization [84, 94] clusters similar things together, it is imperative to be equipped with retrieval capabilities that can bring forward every instance [7]. Approximate Nearest Neighbor Search (ANNS) [40] makes it feasible with efficient indexing [14] and traversal [5, 6] to present the users with the most similar documents/images from the database for a requested query. Widely adopted HNSW [59] ($O(d\log(N))$) is as accurate as exact retrieval ($O(dN)$) at the cost of a graph-based index overhead for RAM and disk [42].
$\rm MRL$} tackles the linear dependence on embedding size, $d$, by learning multifidelity $\rm Matryoshka~Representations$}. Lower-dimensional $\rm Matryoshka~Representations$} are as accurate as independently trained counterparts without the multiple expensive forward passes. $\rm Matryoshka~Representations$} provide an intermediate abstraction between high-dimensional vectors and their efficient ANNS indices through the adaptive embeddings nested within the original representation vector (Section 4). All other aforementioned efficiency techniques are complementary and can be readily applied to the learned $\rm Matryoshka~Representations$} obtained from $\rm MRL$}.
Several works in efficient neural network literature [9, 88, 95] aim at packing neural networks of varying capacity within the same larger network. However, the weights for each progressively smaller network can be different and often require distinct forward passes to isolate the final representations. This is detrimental for adaptive inference due to the need for re-encoding the entire retrieval database with expensive sub-net forward passes of varying capacities. Finally, ordered representations proposed by Rippel et al. [69] use nested dropout in the context of autoencoders to learn nested representations. $\rm MRL$} differentiates itself in formulation by optimizing only for $O(\log(d))$ nesting dimensions instead of $O(d)$. Despite this, $\rm MRL$} diffuses information to intermediate dimensions interpolating between the optimized $\rm Matryoshka~Representation$} sizes accurately (Figure 5); making web-scale feasible.
3 [Uncaptioned image] $\rm Matryoshka~Representation~Learning$}
For $d\in\mathbb{N}$, consider a set $\mathcal{M}\subset[d]$ of representation sizes. For a datapoint $x$ in the input domain $\mathcal{X}$, our goal is to learn a $d$-dimensional representation vector $z\in\mathbb{R}^{d}$. For every $m\in\mathcal{M}$, $\rm Matryoshka~Representation~Learning$} ($\rm MRL$}) enables each of the first $m$ dimensions of the embedding vector, $z_{1:m}\in\mathbb{R}^{m}$ to be independently capable of being a transferable and general purpose representation of the datapoint $x$. We obtain $z$ using a deep neural network $F(\,\cdot\,;\theta_{F})\colon\mathcal{X}\rightarrow\mathbb{R}^{d}$ parameterized by learnable weights $\theta_{F}$, i.e., $z\coloneqq F(x;\theta_{F})$. The multi-granularity is captured through the set of the chosen dimensions $\mathcal{M}$, that contains less than $\log(d)$ elements, i.e., $\lvert\mathcal{M}\rvert\leq\left\lfloor\log(d)\right\rfloor$. The usual set $\mathcal{M}$ consists of consistent halving until the representation size hits a low information bottleneck. We discuss the design choices in Section 4 for each of the representation learning settings.
For the ease of exposition, we present the formulation for fully supervised representation learning via multi-class classification. $\rm Matryoshka~Representation~Learning$} modifies the typical setting to become a multi-scale representation learning problem on the same task. For example, we train ResNet50 [27] on ImageNet-1K [71] which embeds a $224\times 224$ pixel image into a $d=2048$ representation vector and then passed through a linear classifier to make a prediction, $\hat{y}$ among the $L=1000$ labels. For $\rm MRL$}, we choose $\mathcal{M}=\{8,16,\ldots,1024,2048\}$ as the nesting dimensions.
Suppose we are given a labelled dataset $\mathcal{D}=\{(x_{1},y_{1}),\ldots,(x_{N},y_{N})\}$ where $x_{i}\in\mathcal{X}$ is an input point and $y_{i}\in[L]$ is the label of $x_{i}$ for all $i\in[N]$. $\rm MRL$} optimizes the multi-class classification loss for each of the nested dimension $m\in\mathcal{M}$ using standard empirical risk minimization using a separate linear classifier, parameterized by $\mathbf{W}^{(m)}\in\mathbb{R}^{L\times m}$. All the losses are aggregated after scaling with their relative importance $\left(c_{m}\geq 0\right)_{m\in\mathcal{M}}$ respectively. That is, we solve
$\min_{\left\{\mathbf{W}^{(m)}\right\}_{m\in\mathcal{M}},\ \theta_{F}}\frac{1}{N}\sum_{i\in[N]}\sum_{m\in\mathcal{M}}c_{m}\cdot{\cal L}\left(\mathbf{W}^{(m)}\cdot F(x_{i};\theta_{F})_{1:m}\ ;\ y_{i}\right)\ ,$ (1)
where ${\cal L}\colon\mathbb{R}^{L}\times[L]\to\mathbb{R}_{+}$ is the multi-class softmax cross-entropy loss function. This is a standard optimization problem that can be solved using sub-gradient descent methods. We set all the importance scales, $c_{m}=1$ for all $m\in\mathcal{M}$; see Section 5 for ablations. Lastly, despite only optimizing for $O(\log(d))$ nested dimensions, $\rm MRL$} results in accurate representations, that interpolate, for dimensions that fall between the chosen granularity of the representations (Section 4.2).
We call this formulation as $\rm Matryoshka~Representation~Learning$} ($\rm MRL$}). A natural way to make this efficient is through weight-tying across all the linear classifiers, i.e., by defining $\mathbf{W}^{(m)}=\mathbf{W}_{1:m}$ for a set of common weights $\mathbf{W}\in\mathbb{R}^{L\times d}$. This would reduce the memory cost due to the linear classifiers by almost half, which would be crucial in cases of extremely large output spaces [84, 94]. This variant is called Efficient $\rm Matryoshka~Representation~Learning$} ($\rm MRL\text{--}E$}). Refer to Alg 1 and Alg 2 in Appendix A for the building blocks of $\rm Matryoshka~Representation~Learning$} ($\rm MRL$}).
Adaptation to Learning Frameworks.
$\rm MRL$} can be adapted seamlessly to most representation learning frameworks at web-scale with minimal modifications (Section 4.1). For example, $\rm MRL$}’s adaptation to masked language modelling reduces to $\rm MRL\text{--}E$} due to the weight-tying between the input embedding matrix and the linear classifier. For contrastive learning, both in context of vision & vision + language, $\rm MRL$} is applied to both the embeddings that are being contrasted with each other. The presence of normalization on the representation needs to be handled independently for each of the nesting dimension for best results (see Appendix C for more details).
4 Applications
In this section, we discuss $\rm Matryoshka~Representation~Learning$} ($\rm MRL$}) for a diverse set of applications along with an extensive evaluation of the learned multifidelity representations. Further, we showcase the downstream applications of the learned $\rm Matryoshka~Representations$} for flexible large-scale deployment through (a) Adaptive Classification (AC) and (b) Adaptive Retrieval (AR).
Figure 2: ImageNet-1K linear classification accuracy of ResNet50 models. $\rm MRL$} is as accurate as the independently trained FF models for every representation size.
Figure 3: ImageNet-1K 1-NN accuracy of ResNet50 models measuring the representation quality for downstream task. $\rm MRL$} outperforms all the baselines across all representation sizes.
4.1 Representation Learning
We adapt $\rm Matryoshka~Representation~Learning$} ($\rm MRL$}) to various representation learning setups (a) Supervised learning for vision: ResNet50 [27] on ImageNet-1K [71] and ViT-B/16 [22] on JFT-300M [80], (b) Contrastive learning for vision + language: ALIGN model with ViT-B/16 vision encoder and BERT language encoder on ALIGN data [44] and (c) Masked language modelling: BERT [19] on English Wikipedia and BooksCorpus [97]. Please refer to Appendices B and C for details regarding the model architectures, datasets and training specifics.
We do not search for best hyper-parameters for all $\rm MRL$} experiments but use the same hyper-parameters as the independently trained baselines. ResNet50 outputs a $2048$-dimensional representation while ViT-B/16 and BERT-Base output $768$-dimensional embeddings for each data point. We use $\mathcal{M}=\{8,16,32,64,128,256,512,1024,2048\}$ and $\mathcal{M}=\{12,24,48,96,192,384,768\}$ as the explicitly optimized nested dimensions respectively. Lastly, we extensively compare the $\rm MRL$} and $\rm MRL\text{--}E$} models to independently trained low-dimensional (fixed feature) representations (FF), dimensionality reduction (SVD), sub-net method (slimmable networks [95]) and randomly selected features of the highest capacity FF model.
In section 4.2, we evaluate the quality and capacity of the learned representations through linear classification/probe (LP) and 1-nearest neighbour (1-NN) accuracy. Experiments show that $\rm MRL$} models remove the dependence on $|\mathcal{M}|$ resource-intensive independently trained models for the coarse-to-fine representations while being as accurate. Lastly, we show that despite optimizing only for $|\mathcal{M}|$ dimensions, $\rm MRL$} models diffuse the information, in an interpolative fashion, across all the $d$ dimensions providing the finest granularity required for adaptive deployment.
4.2 Classification
[…]
We also evaluate the quality of the representations from training ViT-B/16 on JFT-300M alongside the ViT-B/16 vision encoder of the ALIGN model – two web-scale setups. Due to the expensive nature of these experiments, we only train the highest capacity fixed feature model and choose random features for evaluation in lower-dimensions. Web-scale is a compelling setting for $\rm MRL$} due to its relatively inexpensive training overhead while providing multifidelity representations for downstream tasks. Figure 5, evaluated with 1-NN on ImageNet-1K, shows that all the $\rm MRL$} models for JFT and ALIGN are highly accurate while providing an excellent cost-vs-accuracy trade-off at lower-dimensions. These experiments show that $\rm MRL$} seamlessly scales to large-scale models and web-scale datasets while providing the otherwise prohibitively expensive multi-granularity in the process. We also have similar observations when pretraining BERT; please see Appendix D.2 for more details.
Figure 4: ImageNet-1K 1-NN accuracy for ViT-B/16 models trained on JFT-300M & as part of ALIGN. $\rm MRL$} scales seamlessly to web-scale with minimal training overhead.
Figure 5: Despite optimizing $\rm MRL$} only for $O(\log(d))$ dimensions for ResNet50 and ViT-B/16 models; the accuracy in the intermediate dimensions shows interpolating behaviour.
Our experiments also show that post-hoc compression (SVD), linear probe on random features, and sub-net style slimmable networks drastically lose accuracy compared to $\rm MRL$} as the representation size decreases. Finally, Figure 5 shows that, while $\rm MRL$} explicitly optimizes $O(\log(d))$ nested representations – removing the $O(d)$ dependence [69] –, the coarse-to-fine grained information is interpolated across all $d$ dimensions providing highest flexibility for adaptive deployment.
4.2.1 Adaptive Classification
The flexibility and coarse-to-fine granularity within $\rm Matryoshka~Representations$} allows model cascades [85] for Adaptive Classification (AC) [26]. Unlike standard model cascades [90], $\rm MRL$} does not require multiple expensive neural network forward passes. To perform AC with an $\rm MRL$} trained model, we learn thresholds on the maximum softmax probability [31] for each nested classifier on a holdout validation set. We then use these thresholds to decide when to transition to the higher dimensional representation (e.g $8\to 16\to 32$) of the $\rm MRL$} model. Appendix D.1 discusses the implementation and learning of thresholds for cascades used for adaptive classification in detail.
Figure 7 shows the comparison between cascaded $\rm MRL$} representations ($\rm MRL$}–AC) and independently trained fixed feature (FF) models on ImageNet-1K with ResNet50. We computed the expected representation size for $\rm MRL$}–AC based on the final dimensionality used in the cascade. We observed that $\rm MRL$}–AC was as accurate, $76.30\%$, as a 512-dimensional FF model but required an expected dimensionality of $\sim 37$ while being only $0.8\%$ lower than the 2048-dimensional FF baseline. Note that all $\rm MRL$}–AC models are significantly more accurate than the FF baselines at comparable representation sizes. $\rm MRL$}–AC uses up to $\sim 14\times$ smaller representation size for the same accuracy which affords computational efficiency as the label space grows [84]. Lastly, our results with $\rm MRL$}–AC indicate that instances and classes vary in difficulty which we analyze in Section 5 and Appendix J.
4.3 Retrieval
Nearest neighbour search with learned representations powers a plethora of retrieval and search applications [15, 86, 11, 62]. In this section, we discuss the image retrieval performance of the pretrained ResNet50 models (Section 4.1) on two large-scale datasets ImageNet-1K [71] and ImageNet-4K. ImageNet-1K has a database size of $\sim$1.3M and a query set of 50K samples uniformly spanning 1000 classes. We also introduce ImageNet-4K which has a database size of $\sim$4.2M and query set of $\sim$200K samples uniformly spanning 4202 classes (see Appendix B for details). A single forward pass on ResNet50 costs 4 GFLOPs while exact retrieval costs 2.6 GFLOPs per query for ImageNet-1K. Although retrieval overhead is $40\%$ of the total cost, retrieval cost grows linearly with the size of the database. ImageNet-4K presents a retrieval benchmark where the exact search cost becomes the computational bottleneck ($8.6$ GFLOPs per query). In both these settings, the memory and disk usage are also often bottlenecked by the large databases. However, in most real-world applications exact search, $O(dN)$, is replaced with an approximate nearest neighbor search (ANNS) method like HNSW [59], $O(d\log(N))$, with minimal accuracy drop at the cost of additional memory overhead.
The goal of image retrieval is to find images that belong to the same class as the query using representations obtained from a pretrained model. In this section, we compare retrieval performance using mean Average Precision @ 10 (mAP@$10$) which comprehensively captures the setup of relevant image retrieval at scale. We measure the cost per query using exact search in MFLOPs. All embeddings are unit normalized and retrieved using the L2 distance metric. Lastly, we report an extensive set of metrics spanning mAP@$k$ and P@$k$ for $k=\{10,25,50,100\}$ and real-world wall-clock times for exact search and HNSW. See Appendices E and F for more details.
Figure 6: Adaptive classification on $\rm MRL$} ResNet50 using cascades results in $14\times$ smaller representation size for the same level of accuracy on ImageNet-1K ($\sim 37$ vs $512$ dims for $76.3\%$).
Figure 7: mAP@$10$ for Image Retrieval on ImageNet-1K with ResNet50. $\rm MRL$} consistently produces better retrieval performance over the baselines across all the representation sizes.
Figure 7 compares the mAP@$10$ performance of ResNet50 representations on ImageNet-1K across dimensionalities for $\rm MRL$}, $\rm MRL\text{--}E$}, FF, slimmable networks along with post-hoc compression of vectors using SVD and random feature selection. $\rm Matryoshka~Representations$} are often the most accurate while being up to $3\%$ better than the FF baselines. Similar to classification, post-hoc compression and slimmable network baselines suffer from significant drop-off in retrieval mAP@$10$ with $\leq 256$ dimensions. Appendix E discusses the mAP@$10$ of the same models on ImageNet-4K.
$\rm MRL$} models are capable of performing accurate retrieval at various granularities without the additional expense of multiple model forward passes for the web-scale databases. FF models also generate independent databases which become prohibitively expense to store and switch in between. $\rm Matryoshka~Representations$} enable adaptive retrieval (AR) which alleviates the need to use full-capacity representations, $d=2048$, for all data and downstream tasks. Lastly, all the vector compression techniques [57, 43] used as part of the ANNS pipelines are complimentary to $\rm Matryoshka~Representations$} and can further improve the efficiency-vs-accuracy trade-off.
4.3.1 Adaptive Retrieval
We benchmark $\rm MRL$} in the adaptive retrieval setting (AR) [48]. For a given query image, we obtained a shortlist, $K=200$, of images from the database using a lower-dimensional representation, e.g. $D_{s}=16$ followed by reranking with a higher capacity representation, e.g. $D_{r}=2048$. In real-world scenarios where top ranking performance is the key objective, measured with mAP@$k$ where k covers a limited yet crucial real-estate, AR provides significant compute and memory gains over single-shot retrieval with representations of fixed dimensionality. Finally, the most expensive part of AR, as with any retrieval pipeline, is the nearest neighbour search for shortlisting. For example, even naive re-ranking of 200 images with 2048 dimensions only costs 400 KFLOPs. While we report exact search cost per query for all AR experiments, the shortlisting component of the pipeline can be sped-up using ANNS (HNSW). Appendix I has a detailed discussion on compute cost for exact search, memory overhead of HNSW indices and wall-clock times for both implementations. We note that using HNSW with 32 neighbours for shortlisting does not decrease accuracy during retrieval.
Refer to caption Refer to caption Refer to caption
(a) ImageNet-1K (b) ImageNet-4K
Figure 8: The trade-off between mAP@$10$ vs MFLOPs/Query for Adaptive Retrieval (AR) on ImageNet-1K (left) and ImageNet-4K (right). Every combination of $D_{s}$ & $D_{r}$ falls above the Pareto line (orange dots) of single-shot retrieval with a fixed representation size while having configurations that are as accurate while being up to $14\times$ faster in real-world deployment. Funnel retrieval is almost as accurate as the baseline while alleviating some of the parameter choices of Adaptive Retrieval.
Figure 8 showcases the compute-vs-accuracy trade-off for adaptive retrieval using $\rm Matryoshka~Representations$} compared to single-shot using fixed features with ResNet50 on ImageNet-1K. We observed that all AR settings lied above the Pareto frontier of single-shot retrieval with varying representation sizes. In particular for ImageNet-1K, we show that the AR model with $D_{s}=16$ & $D_{r}=2048$ is as accurate as single-shot retrieval with $d=2048$ while being $\mathbf{\sim 128\times}$ more efficient in theory and $\mathbf{\sim 14\times}$ faster in practice (compared using HNSW on the same hardware). We show similar trends with ImageNet-4K, but note that we require $D_{s}=64$ given the increased difficulty of the dataset. This results in $\sim 32\times$ and $\sim 6\times$ theoretical and in-practice speedups respectively. Lastly, while $K=200$ works well for our adaptive retrieval experiments, we ablated over the shortlist size $k$ in Appendix K.2 and found that the accuracy gains stopped after a point, further strengthening the use-case for $\rm Matryoshka~Representation~Learning$} and adaptive retrieval.
Even with adaptive retrieval, it is hard to determine the choice of $D_{s}$ & $D_{r}$. In order to alleviate this issue to an extent, we propose Funnel Retrieval, a consistent cascade for adaptive retrieval. Funnel thins out the initial shortlist by a repeated re-ranking and shortlisting with a series of increasing capacity representations. Funnel halves the shortlist size and doubles the representation size at every step of re-ranking. For example on ImageNet-1K, a funnel with the shortlist progression of $200\to 100\to 50\to 25\to 10$ with the cascade of $16\to 32\to 64\to 128\to 256\to 2048$ representation sizes within $\rm Matryoshka~Representation$} is as accurate as the single-shot 2048-dim retrieval while being $\sim 128\times$ more efficient theoretically (see Appendix F for more results). All these results showcase the potential of $\rm MRL$} and AR for large-scale multi-stage search systems [15].
5 Further Analysis and Ablations
Robustness.
We evaluate the robustness of the $\rm MRL$} models trained on ImageNet-1K on out-of-domain datasets, ImageNetV2/R/A/Sketch [68, 32, 33, 89], and compare them to the FF baselines. Table 17 in Appendix H demonstrates that $\rm Matryoshka~Representations$} for classification are at least as robust as the original representation while improving the performance on ImageNet-A by $0.6\%$ – a $20\%$ relative improvement. We also study the robustness in the context of retrieval by using ImageNetV2 as the query set for ImageNet-1K database. Table 9 in Appendix E shows that $\rm MRL$} models have more robust retrieval compared to the FF baselines by having up to $3\%$ higher mAP@$10$ performance. This observation also suggests the need for further investigation into robustness using nearest neighbour based classification and retrieval instead of the standard linear probing setup. We also find that the zero-shot robustness of ALIGN-$\rm MRL$} (Table 18 in Appendix H) agrees with the observations made by Wortsman et al. [91]. Lastly, Table 6 in Appendix D.2 shows that $\rm MRL$} also improves the cosine similarity span between positive and random image-text pairs.
Few-shot and Long-tail Learning.
We exhaustively evaluated few-shot learning on $\rm MRL$} models using nearest class mean [74]. Table 15 in Appendix G shows that that representations learned through $\rm MRL$} perform comparably to FF representations across varying shots and number of classes.
$\rm Matryoshka~Representations$} realize a unique pattern while evaluating on FLUID [87], a long-tail sequential learning framework. We observed that $\rm MRL$} provides up to $2\%$ accuracy higher on novel classes in the tail of the distribution, without sacrificing accuracy on other classes (Table 16 in Appendix G). Additionally we find the accuracy between low-dimensional and high-dimensional representations is marginal for pretrain classes. We hypothesize that the higher-dimensional representations are required to differentiate the classes when few training examples of each are known. This results provides further evidence that different tasks require varying capacity based on their difficulty.
Figure 9: Grad-CAM [75] progression of predictions in $\rm MRL$} model across $8,16,32\text{ and }2048$ dimensions. (a) 88-dimensional representation confuses due to presence of other relevant objects (with a larger field of view) in the scene and predicts “shower cap” ; (b) 88-dim model confuses within the same super-class of “boa” ; (c) 88 and $16$-dim models incorrectly focus on the eyes of the doll ("sunglasses") and not the "sweatshirt" which is correctly in focus at higher dimensions; $\rm MRL$} fails gracefully in these scenarios and shows potential use cases of disagreement across dimensions.
Disagreement across Dimensions.
The information packing in $\rm Matryoshka~Representations$} often results in gradual increase of accuracy with increase in capacity. However, we observed that this trend was not ubiquitous and certain instances and classes were more accurate when evaluated with lower-dimensions (Figure 12 in Appendix J). With perfect routing of instances to appropriate dimension, $\rm MRL$} can gain up to $4.6\%$ classification accuracy. At the same time, the low-dimensional models are less accurate either due to confusion within the same superclass [23] of the ImageNet hierarchy or presence of multiple objects of interest. Figure 9 showcases 2 such examples for $8$-dimensional representation. These results along with Appendix J put forward the potential for $\rm MRL$} to be a systematic framework for analyzing the utility and efficiency of information bottlenecks.
Figure 10: 31-way ImageNet-1K superclass classification across representation size for $\rm MRL$} & FF models showing the capture of underlying hierarchy through tight information bottlenecks.
Figure 11: Diverse per-superclass accuracy trends across representation sizes for ResNet50-$\rm MRL$} on ImageNet-1K.
Superclass Accuracy.
As the information bottleneck becomes smaller, the overall accuracy on fine-grained classes decreases rapidly (Figure 3). However, the drop-off is not as significant when evaluated at a superclass level (Table 24 in Appendix J). Figure 11 presents that this phenomenon occurs with both $\rm MRL$} and FF models; $\rm MRL$} is more accurate across dimensions. This shows that tight information bottlenecks while not highly accurate for fine-grained classification, do capture required semantic information for coarser classification that could be leveraged for adaptive routing for retrieval and classification. Mutifidelity of $\rm Matryoshka~Representation$} naturally captures the underlying hierarchy of the class labels with one single model. Lastly, Figure 11 showcases the accuracy trends per superclass with $\rm MRL$}. The utility of additional dimensions in distinguishing a class from others within the same superclass is evident for “garment” which has up to 11% improvement for 8 $\to$ 16 dimensional representation transition. We also observed that superclasses such as “oscine (songbird)” had a clear visual distinction between the object and background and thus predictions using 8 dimensions also led to a good inter-class separability within the superclass.
5.1 Ablations
Table 26 in Appendix K presents that $\rm Matryoshka~Representations$} can be enabled within off-the-shelf pretrained models with inexpensive partial finetuning thus paving a way for ubiquitous adoption of $\rm MRL$}. At the same time, Table 27 in Appendix C indicates that with optimal weighting of the nested losses we could improve accuracy of lower-dimensions representations without accuracy loss. Tables 29 and 29 in Appendix C ablate over the choice of initial granularity and spacing of the granularites. Table 29 reaffirms the design choice to shun extremely low dimensions that have poor classification accuracy as initial granularity for $\rm MRL$} while Table 29 confirms the effectiveness of logarthmic granularity spacing inspired from the behaviour of accuracy saturation across dimensions over uniform. Lastly, Tables 30 and 31 in Appendix K.2 show that the retrieval performance saturates after a certain shortlist dimension and length depending on the complexity of the dataset.
6 Discussion and Conclusions
The results in Section 5.1 reveal interesting weaknesses of $\rm MRL$} that would be logical directions for future work. (1) Optimizing the weightings of the nested losses to obtain a Pareto optimal accuracy-vs-efficiency trade-off – a potential solution could emerge from adaptive loss balancing aspects of anytime neural networks [39]. (2) Using different losses at various fidelities aimed at solving a specific aspect of adaptive deployment – e.g. high recall for $8$-dimension and robustness for $2048$-dimension. (3) Learning a search data-structure, like differentiable k-d tree, on top of $\rm Matryoshka~Representation$} to enable dataset and representation aware retrieval. (4) Finally, the joint optimization of multi-objective $\rm MRL$} combined with end-to-end learnable search data-structure to have data-driven adaptive large-scale retrieval for web-scale search applications.
In conclusion, we presented [Uncaptioned image] $\rm Matryoshka~Representation~Learning$} ($\rm MRL$}), a flexible representation learning approach that encodes information at multiple granularities in a single embedding vector. This enables the $\rm MRL$} to adapt to a downstream task’s statistical complexity as well as the available compute resources. We demonstrate that $\rm MRL$} can be used for large-scale adaptive classification as well as adaptive retrieval. On standard benchmarks, $\rm MRL$} matches the accuracy of the fixed-feature baseline despite using $14\times$ smaller representation size on average. Furthermore, the $\rm Matryoshka~Representation$} based adaptive shortlisting and re-ranking system ensures comparable mAP@$10$ to the baseline while being $128\times$ cheaper in FLOPs and $14\times$ faster in wall-clock time. Finally, most of the efficiency techniques for model inference and vector search are complementary to $\rm MRL$} [Uncaptioned image] further assisting in deployment at the compute-extreme environments.
Acknowledgments
We are grateful to Srinadh Bhojanapalli, Lovish Madaan, Raghav Somani and Ludwig Schmidt for helpful discussions and feedback. Aditya Kusupati also thanks Tom Duerig and Rahul Sukthankar for their support. Part of the paper’s large-scale experimentation is supported through a research GCP credit award from Google Cloud and Google Research. Gantavya Bhatt is supported in part by the CONIX Research Center, one of six centers in JUMP, a Semiconductor Research Corporation (SRC) program sponsored by DARPA. Sham Kakade acknowledges funding from the NSF award CCF-1703574 and ONR N00014-22-1-2377. Ali Farhadi acknowledges funding from the NSF awards IIS 1652052, IIS 17303166, DARPA N66001-19-2-4031, DARPA W911NF-15-1-0543 and gifts from Allen Institute for Artificial Intelligence.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] Our work does not have any additional negative societal impact on top of the existing impact of representation learning. However, a study on the trade-off between representation size and the tendency to encode biases is an interesting future direction along the lines of existing literature [34, 35]. A part of this is already presented in Section 5.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See supplemental material and Appendix A. All the code and public models will be open sourced.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 4 and Appendix C.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] We benchmarked on large-scale datasets like ImageNet-1K, JFT-300M and ALIGN data with models like ResNet and ViT making it extremely expensive to run things multiple times.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix C and Appendix I.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [No] All the non-proprietary datasets and code used are public under MIT, BSD or CC licenses.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We created a new subset of ImageNet-21K for downstream evaluation of retrieval performance at scale. See Section 4.3 and Appendix B
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Code for $\rm Matryoshka~Representation~Learning$} [Uncaptioned image] ($\rm MRL$})
[…]
Algorithm 1 Pytorch code for $\rm Matryoshka$} Cross-Entropy Loss
class Matryoshka_CE_Loss($n$.Module):
def __init__(self, relative_importance, **kwargs):
super(Matryoshka_CE_Loss, self).__init__()
self.criterion = $n$.CrossEntropyLoss(**kwargs)
self.relative_importance = relative_importance # usually set to all ones
def forward(self, output, target):
loss=0
for i in range(len(output)):
loss+= self.relative_importance[i] * self.criterion(output[i], target)
return loss
Algorithm 2 Pytorch code for $\rm MRL$} Linear Layer
class MRL_Linear_Layer($n$.Module):
def __init__(self, nesting_list: List, num_classes=1000, efficient=False, **kwargs):
super(MRL_Linear_Layer, self).__init__()
self.nesting_list=nesting_list # set of m in M (Eq. 1)
self.num_classes=num_classes
self.is_efficient=efficient # flag for MRL-E
if not is_efficient:
for i, num_feat in enumerate(self.nesting_list):
[…]
else:
setattr(self, "nesting_classifier_0", $n$.Linear(self.nesting_list[-1], self.num_classes, **kwargs)) # Instantiating one $n$.Linear layer for MRL-E
def forward(self, x):
nesting_logits = ()
for i, num_feat in enumerate(self.nesting_list):
if(self.is_efficient):
efficient_logit = torch.matmul(x[:, :num_feat], (self.nesting_classifier_0.weight[:, :num_feat]).t())
else:
[…]
if(self.is_efficient):
nesting_logits.append(efficient_logit)
return nesting_logits
Appendix B Datasets
ImageNet-1K [71] contains 1,281,167 labeled train images, and 50,000 labelled validation images across 1,000 classes. The images were transformed with standard procedures detailed by FFCV [54].
ImageNet-4K dataset was constructed by selecting 4,202 classes, non-overlapping with ImageNet-1K, from ImageNet-21K [16] with 1,050 or more examples. The train set contains 1,000 examples and the query/validation set contains 50 examples per class totalling to $\sim$4.2M and $\sim$200K respectively. We will release the list of images curated together to construct ImageNet-4K.
JFT-300M [80] is a large-scale multi-label dataset with 300M images labelled across 18,291 categories.
ALIGN [44] utilizes a large scale noisy image-text dataset containing 1.8B image-text pairs.
ImageNet Robustness Datasets
We experimented on the following datasets to examine the robustness of $\rm MRL$} models:
ImageNetV2 [68] is a collection of 10K images sampled a decade after the original construction of ImageNet [16]. ImageNetV2 contains 10 examples each from the 1,000 classes of ImageNet-1K.
ImageNet-A [33] contains 7.5K real-world adversarially filtered images from 200 ImageNet-1K classes.
ImageNet-R [32] contains 30K artistic image renditions for 200 of the original ImageNet-1K classes.
ImageNet-Sketch [89] contains 50K sketches, evenly distributed over all 1,000 ImageNet-1K classes.
ObjectNet [2] contains 50K images across 313 object classes, each containing $\sim$160 images each.
Appendix C $\rm Matryoshka~Representation~Learning$} Model Training
[…]
  • $\rm MRL$}: ResNet50 model with the fc layer replaced by MRL_Linear_Layer(efficient=False)
  • $\rm MRL\text{--}E$}: ResNet50 model with the fc layer replaced by MRL_Linear_Layer(efficient=True)
  • FF–k: ResNet50 model with the fc layer replaced by torch.nn.Linear(k, num_classes), where k $\in[8,16,32,64,128,256,512,1024,2048]$. We will henceforth refer to these models as simply FF, with the k value denoting representation size.
[…]
Our code (Appendix A) makes minimal modifications to the training pipeline provided by FFCV to learn $\rm Matryoshka~Representations$}.
We trained ViT-B/16 models for JFT-300M on a 8x8 cloud TPU pod [47] using Tensorflow [1] with a batchsize of 128 and trained for 300K steps. Similarly, ALIGN models were trained using Tensorflow on 8x8 cloud TPU pod for 1M steps with a batchsize of 64 per TPU. Both these models were trained with adafactor optimizer [76] with a linear learning rate decay starting at 1e-3.
Lastly, we trained a BERT-Base model on English Wikipedia and BookCorpus. We trained our models in Tensorflow using a 4x4 cloud TPU pod with a total batchsize of 1024. We used AdamW [58] optimizer with a linear learning rate decay starting at 1e-4 and trained for 450K steps.
In each configuration/case, if the final representation was normalized in the FF implementation, $\rm MRL$} models adopted the same for each nested dimension for a fair comparison.
Appendix D Classification Results
Table 1: Top-1 classification accuracy (%) for ResNet50 $\rm MRL$} and baseline models on ImageNet-1K.
Rep. Size Rand. LP SVD $F$ Slim. Net $\rm MRL$} $\rm MRL\text{--}E$}
8  4.56  2.34 65.29  0.42 66.63 56.66
16 11.29  7.17 72.85  0.96 73.53 71.94
32 27.21 20.46 74.60  2.27 75.03 74.48
64 49.47 48.10 75.27  5.59 75.82 75.35
128 65.70 67.24 75.29 14.15 76.30 75.80
256 72.43 74.59 75.71 38.42 76.47 76.22
512 74.94 76.78 76.18 69.80 76.65 76.36
1024 76.10 76.87 76.63 74.61 76.76 76.48
2048 76.87 76.87 76.26 76.80 76.51
[…]
  • FF: We utilize the FF-k models described in Appendix C for $k\in\{8,...2048\}$.
  • SVD: We performed a low rank approximation of the 1000-way classification layer of FF-2048, with rank = 1000.
  • Rand. LP: We compared against a linear classifier fit on randomly selected features [28].
  • Slim. Net: We take pretrained slimmable neural networks [95] which are trained with a flexible width backbone (25%, 50%, 75% and full width). For each representation size, we consider the first $k$ dimensions for classification. Note that training of slimmable neural networks becomes unstable when trained below 25% width due to the hardness in optimization and low complexity of the model.
At lower dimensions ( $d\leq 128$), $\rm MRL$} outperforms all baselines significantly, which indicates that pretrained models lack the multifidelity of $\rm Matryoshka~Representations$} and are incapable of fitting an accurate linear classifier at low representation sizes.
We compared the performance of $\rm MRL$} models at various representation sizes via 1-nearest neighbors (1-NN) image classification accuracy on ImageNet-1K in Table 2 and Figure 3. We provide detailed information regarding the k-NN search pipeline in Appendix E. We compared against a baseline of attempting to enforce nesting to a FF-2048 model by 1) Random Feature Selection (Rand. FS): considering the first $m$ dimensions of FF-2048 for NN lookup, and 2) FF+SVD: performing SVD on the FF-2048 representations at the specified representation size, 3) FF+JL: performing random projection according to the Johnson-Lindenstrauss lemma [46] on the FF-2048 representations at the specified representation size. We also compared against the 1-NN accuracy of slimmable neural nets [95] as an additional baseline. We observed these baseline models to perform very poorly at lower dimensions, as they were not explicitly trained to learn $\rm Matryoshka~Representations$}.
Table 2: 1-NN accuracy (%) on ImageNet-1K for various ResNet50 models.
Rep. Size Rand. FS SVD JL $F$ Slimmable $\rm MRL$} $\rm MRL\text{--}E$}
8  2.36 19.14 0.11 58.93  1.00 62.19 57.45
16 12.06 46.02 0.09 66.77  5.12 67.91 67.05
32 32.91 60.78 0.06 68.84 16.95 69.46 68.6
64 49.91 67.04 0.05 69.41 35.60 70.17 69.61
128 60.91 69.63 0.06 69.35 51.16 70.52 70.12
256 65.75 70.67 0.04 69.72 60.61 70.62 70.36
512 68.77 71.06 0.03 70.18 65.82 70.82 70.74
1024 70.41 71.22 - 70.34 67.19 70.89 71.07
2048 71.19 71.21 - 71.19 66.10 70.97 71.21
D.1 Adaptive Classification ($\rm MRL$}–AC)
Table 3: Threshold-based adaptive classification performance of ResNet50 $\rm MRL$} on a 40K sized held-out subset of the ImageNet-1K validation set. Results are averaged over 30 random held-out subsets.
Expected Rep. Size Accuracy
 13.43 $\pm$  0.81 73.79 $\pm$ 0.10
 18.32 $\pm$  1.36 75.25 $\pm$ 0.11
 25.87 $\pm$  2.41 76.05 $\pm$ 0.15
 36.26 $\pm$  4.78 76.28 $\pm$ 0.16
 48.00 $\pm$  8.24 76.43 $\pm$ 0.18
 64.39 $\pm$  12.55 76.53 $\pm$ 0.19
 90.22 $\pm$ 20.88 76.55 $\pm$ 0.20
 118.85 $\pm$ 33.37 76.56 $\pm$ 0.20
In an attempt to use the smallest representation that works well for classification for every image in the ImageNet-1K validation set, we learned a policy to increase the representation size from $m_{i}$ to $m_{i+1}$ using a 10K sized subset of the ImageNet-1K validation set. This policy is based on whether the prediction confidence $p_{i}$ using representation size $m_{i}$ exceeds a learned threshold $t_{i}^{\ast}$. If $p_{i}\geq t_{i}^{\ast}$, we used predictions from representation size $m_{i}$ otherwise, we increased to representation size $m_{i+1}$. To learn the optimal threshold $t_{i}^{\ast}$, we performed a grid search between 0 and 1 (100 samples). For each threshold $t_{k}$, we computed the classification accuracy over our 10K image subset. We set $t_{i}^{\ast}$ equal to the smallest threshold $t_{k}$ that gave the best accuracy. We use this procedure to obtain thresholds for successive models, i.e., $\{t_{j}^{\ast}\mid j\in\{8,16,32,64,\ldots,2048\}\}$. To improve reliability of threshold based greedy policy, we use test time augmentation which has been used successfully in the past [77].
For inference, we used the remaining held-out 40K samples from the ImageNet-1K validation set. We began with smallest sized representation ($m=8$) and compared the computed prediction confidence $p_{8}$ to learned optimal threshold $t_{8}^{\ast}$. If $p_{8}\leq t_{8}^{\ast}$, then we increased $m=16$, and repeated this procedure until $m=d=2048$. To compute the expected dimensions, we performed early stopping at $m=\{16,32,64,\ldots 2048\}$ and computed the expectation using the distribution of representation sizes. As shown in Table 3 and Figure 7, we observed that in expectation, we only needed a $\sim 37$ sized representation to achieve $76.3\%$ classification accuracy on ImageNet-1K, which was roughly $14\times$ smaller than the FF–512 baseline. Even if we computed the expectation as a weighted average over the cumulative sum of representation sizes $\{8,24,56,\ldots\}$, due to the nature of multiple linear heads for $\rm MRL$}, we ended up with an expected size of $62$ that still provided a roughly $8.2\times$ efficient representation than the FF–512 baseline. However, $\rm MRL\text{--}E$} alleviates this extra compute with a minimal drop in accuracy.
D.2 JFT, ALIGN and BERT
[…]
Table 4: ViT-B/16 and ViT-B/16-$\rm MRL$} top-1 and top-5 k-NN accuracy (%) for ALIGN and JFT. Top-1 entries where $\rm MRL\text{--}E$} and $\rm MRL$} outperform baselines are bolded for both ALIGN and JFT-ViT.
Rep. Size
ALIGN ALIGN-$\rm MRL$} JFT-ViT ${\rm MRL} JFT-ViT$-$\rm MRL\text{--}E$}
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
12 11.90 28.05 43.57 67.36 27.07 48.57 53.61 75.30 51.54 73.94
24 33.35 55.58 56.44 78.19 48.64 70.20 62.80 81.51 62.40 81.36
48 51.32 73.15 62.33 82.30 63.58 81.80 67.24 84.37 66.89 83.80
96 61.82 81.97 65.72 84.61 68.56 85.13 69.74 85.86 68.80 85.13
192 66.71 85.27 67.00 85.36 71.32 86.21 71.34 86.62 70.41 86.01
384 67.65 85.70 67.70 85.73 71.67 86.98 71.73 87.08 71.18 86.46
768 68.00 86.10 67.85 85.85 72.10 87.20 71.85 86.92 71.31 86.62
Table 5: Examining top-1 and top-5 k-NN accuracy (%) at interpolated hidden dimensions for ALIGN and JFT. This indicates that $\rm MRL$} is able to scale classification accuracy as hidden dimensions increase even at dimensions that were not explicitly considered during training.
Interpolated
Rep. Size
ALIGN-$\rm MRL$} JFT-ViT-$\rm MRL$}
Top-1 Top-5 Top-1 Top-5
16 49.06 72.26 58.35 78.55
32 58.64 79.96 64.98 82.89
64 63.90 83.39 68.19 84.85
128 66.63 85.00 70.35 86.24
256 67.10 85.30 71.57 86.77
512 67.64 85.72 71.55 86.67
Table 6: Cosine similarity between embeddings
Avg. Cosine Similarity ALIGN ALIGN-$\rm MRL$}
Positive Text to Image 0.27 0.49
Random Text to Image 8e-3 -4e-03
Random Image to Image 0.10 0.08
Random Text to Text 0.22 0.07
[…]
Table 7: Masked Language Modelling (MLM) accuracy(%) of FF and $\rm MRL$} models on the validation set.
Rep. Size BERT-$F$ BERT-$\rm MRL$}
12 60.12 59.92
24 62.49 62.05
48 63.85 63.40
96 64.32 64.15
192 64.70 64.58
384 65.03 64.81
768 65.54 65.00
Appendix E Image Retrieval
We evaluated the strength of $\rm Matryoshka~Representations$} via image retrieval on ImageNet-1K (the training distribution), as well as on out-of-domain datasets ImageNetV2 and ImageNet-4K for all $\rm MRL$} ResNet50 models. We generated the database and query sets, containing $N$ and $Q$ samples respectively, with a standard PyTorch [63] forward pass on each dataset. We specify the representation size at which we retrieve a shortlist of k-nearest neighbors (k-NN) by $D_{s}$. The database is a thus a [$N$, $D_{s}$] array, the query set is a [$Q$, $D_{s}$] array, and the neighbors set is a [$Q$, k] array. For metrics, we utilized corrected mean average precision (mAP@k) [53] and precision (P@k): $P@k=\dfrac{correct\_pred}{k}$ where correct_predcorrect\_pred is the average number of retrieved NN with the correct label over the entire query set using a shortlist of length $k$.
We performed retrieval with FAISS [45], a library for efficient similarity search. To obtain a shortlist of k-NN, we built an index to search the database. We performed an exhaustive NN search with the L2 distance metric with faiss.IndexFlatL2, as well as an approximate NN search (ANNS) via HNSW [45] with faiss.IndexHNSWFlat. We used HNSW with $M=32$ unless otherwise mentioned, and henceforth referred to as HNSW32. The exact search index was moved to the GPU for fast k-NN search computation, whereas the HNSW index was kept on the CPU as it currently lacks GPU support. We show the wall clock times for building the index as well as the index size in Table 20. We observed exact search to have a smaller index size which was faster to build when compared to HNSW, which trades off a larger index footprint for fast NN search (discussed in more detail in Appendix K). The database and query vectors are normalized with faiss.normalize_L2 before building the index and performing search.
Table 8: Retrieve a shortlist of 200-NN with $D_{s}$ sized representations on ImageNet-1K via exact search with L2 distance metric. Top-1 and mAP@10 entries (%) where $\rm MRL\text{--}E$} and $\rm MRL$} outperform $F$  at their respective representation sizes are bolded.
Model $D_{s}$ MFlops Top-1 Top-5 Top-10 mAP@10 mAP@25 mAP@50 mAP@100 P@10 P@25 P@50 P@100
$F$ 8 10 58.93 75.76 80.25 53.42 52.29 51.84 51.57 59.32 59.28 59.25 59.21
16 20 66.77 80.88 84.40 61.63 60.51 59.98 59.62 66.76 66.58 66.43 66.27
32 41 68.84 82.58 86.14 63.35 62.08 61.36 60.76 68.43 68.13 67.83 67.48
64 82 69.41 83.56 87.33 63.26 61.64 60.63 59.67 68.49 67.91 67.38 66.74
128 164 69.35 84.23 88.24 62.30 60.16 58.73 57.29 67.84 66.83 65.96 64.92
256 328 69.72 84.71 88.54 61.47 58.85 57.02 55.13 67.19 65.82 64.64 63.24
512 656 70.18 85.04 88.91 61.37 58.41 56.26 53.98 67.12 65.49 64.07 62.35
1024 1312 70.34 85.38 89.19 61.13 57.87 55.47 52.90 66.93 65.08 63.43 61.45
2048 2624 71.19 85.66 89.17 62.90 60.06 57.99 55.76 68.46 66.9 65.52 63.83
$\rm MRL\text{--}E$} 8 10 57.39 74.18 79.16 51.80 50.41 49.60 48.86 57.50 57.16 56.81 56.36
16 20 67.08 81.38 85.15 61.60 60.36 59.66 59.04 66.79 66.53 66.24 65.87
32 41 68.62 82.92 86.44 63.34 61.97 61.14 60.39 68.49 68.06 67.65 67.17
64 82 69.56 83.49 86.85 63.84 62.33 61.43 60.57 68.93 68.4 67.96 67.38
128 164 70.13 83.63 87.07 64.15 62.58 61.61 60.70 69.19 68.62 68.11 67.50
256 328 70.39 83.8 87.28 64.35 62.76 61.76 60.82 69.36 68.79 68.26 67.63
512 656 70.74 83.91 87.33 64.69 63.05 62.06 61.14 69.63 69.00 68.50 67.88
1024 1312 71.05 84.13 87.46 64.85 63.22 62.19 61.26 69.78 69.16 68.60 67.99
2048 2624 71.17 84.27 87.67 64.99 63.33 62.29 61.33 69.90 69.24 68.68 68.05
$\rm MRL\text{--}E$}
Interpolated
12 15 64.25 79.21 83.29 58.83 57.50 56.71 56.02 64.10 63.78 63.42 63.02
24 31 68.28 82.31 85.89 62.75 61.41 60.62 59.92 67.89 67.49 67.11 66.69
48 61 69.20 83.15 86.67 63.58 62.12 61.23 60.42 68.71 68.19 67.75 67.22
96 123 70.05 83.63 87.11 64.04 62.46 61.52 60.63 69.10 68.51 68.04 67.45
192 246 70.36 83.72 87.21 64.26 62.65 61.65 60.72 69.26 68.67 68.15 67.53
384 492 70.54 83.88 87.28 64.55 62.94 61.93 61.01 69.51 68.92 68.40 67.78
768 984 70.96 84.05 87.44 64.79 63.15 62.15 61.22 69.72 69.10 68.56 67.95
1536 1968 71.19 84.17 87.57 64.94 63.29 62.26 61.32 69.85 69.21 68.66 68.04
$\rm MRL$} 8 10 62.19 77.05 81.34 56.74 55.47 54.76 54.12 62.06 61.81 61.54 61.17
16 20 67.91 81.44 85.00 62.94 61.79 61.16 60.64 67.93 67.71 67.48 67.20
32 41 69.46 83.01 86.30 64.21 62.96 62.22 61.58 69.18 68.87 68.54 68.17
64 82 70.17 83.53 86.95 64.69 63.33 62.53 61.80 69.67 69.25 68.89 68.42
128 164 70.52 83.98 87.25 64.94 63.50 62.63 61.83 69.93 69.44 69.02 68.50
256 328 70.62 84.17 87.38 65.04 63.56 62.66 61.81 70.02 69.52 69.07 68.50
512 656 70.82 84.31 87.55 65.14 63.57 62.62 61.73 70.12 69.53 69.04 68.45
1024 1312 70.89 84.44 87.68 65.16 63.58 62.60 61.68 70.14 69.54 69.01 68.41
2048 2624 70.97 84.41 87.74 65.20 63.57 62.56 61.60 70.18 69.52 68.98 68.35
$\rm MRL$}
Interpolated
12 15 65.89 80.04 83.68 60.84 59.66 58.98 58.37 65.94 65.72 65.45 65.08
24 31 68.76 82.48 85.87 63.64 62.42 61.74 61.13 68.64 68.35 68.07 67.71
48 61 69.96 83.40 86.65 64.58 63.2 62.42 61.72 69.53 69.10 68.75 68.32
96 123 70.40 83.83 87.04 64.86 63.46 62.62 61.84 69.82 69.38 68.98 68.48
192 246 70.64 84.09 87.37 65.00 63.53 62.66 61.83 69.98 69.49 69.05 68.50
384 492 70.69 84.25 87.41 65.09 63.56 62.64 61.76 70.05 69.51 69.04 68.46
768 984 70.84 84.40 87.63 65.16 63.59 62.62 61.71 70.14 69.55 69.03 68.44
1536 1968 70.88 84.39 87.71 65.18 63.59 62.58 61.64 70.16 69.54 68.99 68.38
Table 9: Retrieve a shortlist of 200-NN with $D_{s}$ sized representations on ImageNetV2 via exact search with L2 distance metric. Top-1 and mAP@10 entries (%) where $\rm MRL\text{--}E$} outperforms FF  are bolded. $\rm MRL$} outperforms FF at all $D_{s}$ and is thus not bolded.
Config $D_{s}$ MFLOPs Top-1 Top-5 Top-10 mAP@10 mAP@25 mAP@50 mAP@100 P@10 P@25 P@50 P@100
$F$ 8   10 48.79 64.70 69.72 43.04 41.89 41.42 41.17 48.43 48.27 48.25 48.19
16   20 55.08 69.50 74.08 49.63 48.53 48.06 47.75 54.76 54.64 54.53 54.39
32   41 56.69 71.10 76.47 51.11 49.85 49.17 48.65 56.23 55.96 55.71 55.42
64   82 57.37 72.71 77.48 51.28 49.75 48.85 47.99 56.65 56.14 55.71 55.15
128  164 57.17 73.31 78.64 50.07 48.09 46.79 45.58 55.75 54.89 54.12 53.28
256  328 57.09 74.04 79.24 49.11 46.66 44.99 43.35 55.02 53.77 52.74 51.53
512  656 57.12 73.91 79.32 48.95 46.25 44.37 42.42 54.88 53.49 52.29 50.83
1024 1312 57.53 74.17 79.55 48.27 45.41 43.36 41.26 54.31 52.84 51.49 49.87
2048 2624 57.84 74.59 79.45 49.99 47.47 45.66 43.87 55.89 54.63 53.45 52.12
$\rm MRL\text{--}E$} 8   10 47.05 62.53 67.60 40.79 39.47 38.78 38.16 46.03 45.77 45.54 45.17
16   20 55.73 70.54 74.86 49.86 48.57 47.84 47.26 54.97 54.71 54.44 54.10
32   41 57.33 71.61 76.64 51.26 49.92 49.09 48.42 56.46 56.11 55.70 55.30
64   82 57.90 72.55 77.44 51.89 50.29 49.34 48.53 57.06 56.45 55.97 55.43
128  164 57.73 72.79 77.28 52.02 50.38 49.49 48.62 57.13 56.58 56.15 55.58
256  328 58.22 72.77 77.67 52.16 50.61 49.67 48.81 57.30 56.79 56.33 55.77
512  656 58.46 73.00 77.88 52.52 50.97 50.02 49.16 57.65 57.10 56.64 56.08
1024 1312 58.71 73.29 78.00 52.70 51.13 50.17 49.30 57.83 57.26 56.77 56.20
2048 2624 58.86 73.17 78.00 52.88 51.25 50.26 49.36 57.95 57.35 56.85 56.25
$\rm MRL$} 8   10 50.41 65.56 70.27 45.51 44.38 43.71 43.17 50.55 50.44 50.17 49.91
16   20 56.64 70.19 74.61 50.98 49.76 49.16 48.69 55.90 55.66 55.52 55.29
32   41 57.96 71.88 76.41 52.06 50.78 50.09 49.54 57.18 56.83 56.57 56.27
64   82 58.94 72.74 77.17 52.65 51.24 50.44 49.76 57.72 57.29 56.94 56.52
128  164 59.13 73.07 77.49 52.94 51.42 50.53 49.74 58.00 57.47 57.05 56.55
256  328 59.18 73.64 77.75 52.96 51.45 50.52 49.70 58.01 57.53 57.06 56.54
512  656 59.40 73.85 77.97 53.01 51.39 50.46 49.61 58.11 57.49 57.04 56.48
1024 1312 59.11 73.77 77.92 52.98 51.37 50.40 49.54 58.13 57.51 57.00 56.45
2048 2624 59.63 73.84 77.97 52.96 51.34 50.34 49.44 58.07 57.48 56.95 56.36
Table 10: Retrieve a shortlist of 200-NN with $D_{s}$  sized representations on ImageNet-4K  via exact search with L2 distance metric. $\rm MRL\text{--}E$}  and FF  models are omitted for clarity and compute/inference time costs. All entries are in %.
Config $D_{s}$ MFLOPs Top-1 Top-5 Top-10 mAP@10 mAP@25 mAP@50 mAP@100 P@10 P@25 P@50 P@100
$\rm MRL$} 8   34 10.60 26.23 35.57  5.32  4.29  3.76 3.36  9.13  8.77  8.46  8.13
16   67 16.74 36.91 47.28  8.64  6.83  5.84 5.05 13.82 12.79 12.04 13.27
32  134 21.54 43.75 54.11 11.36  8.88  7.47 6.31 17.25 15.67 14.47 13.27
64  269 25.00 47.97 58.25 13.38 10.40  8.67 7.23 19.68 17.64 16.14 14.65
128  538 27.27 50.35 60.47 14.77 11.47  9.53 7.91 21.25 18.95 17.26 15.59
256 1076 28.53 51.95 61.90 15.66 12.19 10.12 8.38 22.28 19.81 18.01 16.22
512 2151 29.46 53.03 62.81 16.29 12.70 10.55 8.72 22.96 20.42 18.54 16.68
1024 4303 30.23 53.72 63.45 16.76 13.08 10.86 8.97 23.48 20.88 18.93 17.00
2048 8606 30.87 54.32 64.02 17.20 13.43 11.14 9.19 23.97 21.28 19.28 17.30
$\rm MRL$}-
Interpolated
12   50 14.04 32.56 42.71  7.16  5.70  4.92 4.32 11.81 11.08 10.52  9.94
24  101 19.49 40.82 51.26 10.17  7.98  6.75 5.75 15.76 14.43 13.42 12.40
48  202 23.51 46.23 56.56 12.49  9.72  8.13 6.81 18.62 16.75 15.39 14.04
96  403 26.25 49.32 59.48 14.15 11.00  9.15 7.61 20.55 18.36 16.78 15.17
192  807 27.94 51.32 61.32 15.29 11.89  9.88 8.18 21.86 19.46 17.71 15.96
384 1614 29.03 52.53 62.45 15.99 12.46 10.35 8.56 22.64 20.14 18.29 16.47
768 3227 29.87 53.36 63.13 16.54 12.90 10.71 8.85 23.23 20.67 18.75 16.85
1536 6454 30.52 54.02 63.79 16.99 13.27 11.01 9.08 23.73 21.09 19.12 17.16
Retrieval performance on ImageNet-1K, i.e. the training distribution, is shown in Table  8. $\rm MRL$} outperforms FF models for nearly all representation size for both top-1 and mAP@10, and especially at low representation size ($D_{s}$ $\leq 32$). $\rm MRL\text{--}E$} loses out to FF significantly only at $D_{s}$ =8=8. This indicates that training ResNet50 models via the $\rm MRL$} training paradigm improves retrieval at low representation size over models explicitly trained at those representation size (FF-820488...2048).
We carried out all retrieval experiments at $D_{s}$ $\in\{8,16,32,64,128,256,512,1024,2048\}$, as these were the representation sizes which were a part of the nesting_list at which losses were added during training, as seen in Algorithm 1, Appendix A. To examine whether $\rm MRL$} is able to learn $\rm Matryoshka~Representations$} at dimensions in between the representation size for which it was trained, we also tabulate the performance of $\rm MRL$} at interpolated $D_{s}$ $\in\{12,24,48,96,192,384,768,1536\}$ as $\rm MRL$}–Interpolated and $\rm MRL\text{--}E$}–Interpolated (see Table 8). We observed that performance scaled nearly monotonically between the original representation size and the interpolated representation size as we increase $D_{s}$, which demonstrates that $\rm MRL$} is able to learn $\rm Matryoshka~Representations$} at nearly all representation size $m\in[8,2048]$ despite optimizing only for $|\mathcal{M}|$ nested representation sizes.
We examined the robustness of $\rm MRL$} for retrieval on out-of-domain datasets ImageNetV2 and ImageNet-4K, as shown in Table 9 and Table 10 respectively. On ImageNetV2, we observed that $\rm MRL$} outperformed FF at all $D_{s}$ on top-1 Accuracy and mAP@10, and $\rm MRL\text{--}E$} outperformed FF  at all $D_{s}$  except $D_{s}$ =8=8. This demonstrates the robustness of the learned $\rm Matryoshka~Representations$} for out-of-domain image retrieval.
Appendix F Adaptive Retrieval
[…]
Table 11: Retrieve a shortlist of k-NN with $D_{s}$ sized representations on ImageNet-1K with $\rm MRL$} representations, and then re-order the neighbors shortlist with L2 distances using $D_{r}$ sized representations. Top-1 and mAP@10 entries (%) that are within $0.1\%$ of the maximum value achievable without reranking on $\rm MRL$} representations, as seen in Table 8, are bolded.
Shortlist Length = 200
$D_{s}$ $D_{r}$ MFLOPs Top-1 mAP@10 mAP@25 mAP@50 mAP@100 P@10 P@25 P@50 P@100
8   16   10 68.21 63.35 62.25 61.70 61.19 68.32 68.14 67.96 67.65
 32 69.42 64.12 62.81 62.03 61.32 69.04 68.63 68.22 67.71
 64 70.05 64.46 63.03 62.14 61.29 69.37 68.83 68.32 67.66
 128 70.34 64.68 63.16 62.21 61.27 69.59 68.96 68.38 67.65
 256 70.40 64.77 63.21 62.23 61.26 69.66 69.02 68.41 67.65
 512 70.60 64.86 63.22 62.21 61.22 69.74 69.02 68.39 67.62
1024 70.71 64.88 63.23 62.20 61.20 69.76 69.01 68.39 67.60
2048 70.81 64.90 63.22 62.17 61.16 69.77 68.99 68.36 67.57
16  32   21 69.47 64.27 63.04 62.36 61.75 69.21 68.90 68.58 68.12
 64 70.16 64.74 63.42 62.66 61.94 69.66 69.22 68.81 68.22
 128 70.52 65.00 63.60 62.77 61.98 69.91 69.36 68.89 68.24
 256 70.55 65.10 63.67 62.82 62.01 69.98 69.43 68.92 68.25
 512 70.74 65.21 63.70 62.83 62.00 70.08 69.43 68.92 68.24
1024 70.83 65.23 63.72 62.83 61.99 70.08 69.45 68.92 68.23
2048 70.90 65.27 63.73 62.82 61.97 70.10 69.44 68.90 68.21
32  64   41 70.16 64.69 63.35 62.57 61.93 69.68 69.26 68.92 68.51
 128 70.52 64.97 63.54 62.73 62.04 69.95 69.47 69.06 68.59
 256 70.63 65.07 63.63 62.79 62.07 70.04 69.55 69.12 68.61
 512 70.82 65.17 63.66 62.80 62.06 70.11 69.57 69.12 68.60
1024 70.89 65.20 63.68 62.80 62.04 70.15 69.59 69.12 68.59
2048 70.97 65.24 63.70 62.79 62.02 70.19 69.59 69.10 68.56
64  128   82 70.51 64.94 63.50 62.64 61.88 69.94 69.44 69.02 68.54
 256 70.63 65.04 63.57 62.69 61.91 70.02 69.52 69.08 68.57
 512 70.83 65.14 63.59 62.67 61.87 70.12 69.54 69.06 68.54
1024 70.89 65.16 63.59 62.65 61.85 70.15 69.54 69.05 68.52
2048 70.97 65.20 63.59 62.63 61.82 70.18 69.53 69.03 68.49
128  256  164 70.63 65.04 63.56 62.66 61.82 70.02 69.52 69.07 68.51
 512 70.82 65.14 63.58 62.63 61.77 70.11 69.54 69.04 68.47
1024 70.89 65.16 63.58 62.60 61.73 70.14 69.54 69.02 68.45
2048 70.97 65.20 63.57 62.57 61.68 70.18 69.52 68.99 68.41
256  512  328 70.82 65.14 63.57 62.62 61.74 70.12 69.53 69.04 68.45
1024 70.88 65.16 63.58 62.60 61.69 70.14 69.54 69.01 68.41
2048 70.97 65.20 63.56 62.56 61.62 70.18 69.52 68.98 68.37
512 1024  656 70.90 65.16 63.58 62.60 61.68 70.14 69.54 69.01 68.41
2048 70.98 65.20 63.57 62.56 61.60 70.18 69.52 68.98 68.35
1024 2048 1312 70.97 65.20 63.57 62.56 61.60 70.18 69.52 68.98 68.35
Table 12: Retrieve a shortlist of k-NN with $D_{s}$ sized representations on ImageNet-4K with $\rm MRL$} representations, and then re-order the neighbors shortlist with L2 distances using $D_{r}$ sized representations. Top-1 and mAP@10 entries (%) that are within $0.1\%$ of the maximum value achievable without reranking on $\rm MRL$} representations, as seen in Table 10, are bolded.
Shortlist Length = 200
$D_{s}$ $D_{r}$ MFLOPs Top-1 mAP@10 mAP@25 mAP@50 mAP@100 P@10 P@25 P@50 P@100
8   16   34 16.84  8.70  6.88 5.88 5.08 13.86 12.80 11.98 11.10
 32 20.73 10.66  8.19 6.77 5.61 16.18 14.39 13.02 11.61
 64 23.11 11.91  9.03 7.36 6.00 17.56 15.34 13.67 11.99
 128 24.63 12.71  9.59 7.76 6.25 18.42 15.94 14.08 12.22
 256 25.5 13.24  9.96 8.03 6.42 19.00 16.35 14.36 12.37
 512 26.07 13.59 10.21 8.20 6.53 19.37 16.62 14.54 12.46
1024 26.52 13.85 10.40 8.34 6.61 19.65 16.80 14.68 12.53
2048 26.94 14.11 10.57 8.45 6.68 19.92 16.98 14.79 12.58
16  32   67 21.44 11.24  8.72 7.26 6.02 17.02 15.30 13.92 12.41
 64 24.36 12.78  9.75 7.96 6.43 18.72 16.41 14.63 12.74
 128 26.08 13.70 10.39 8.39 6.69 19.68 17.07 15.05 12.94
 256 26.99 14.27 10.79 8.67 6.85 20.27 17.48 15.31 13.07
 512 27.60 14.66 11.06 8.86 6.97 20.67 17.75 15.50 13.16
1024 28.12 14.94 11.26 8.99 7.05 20.96 17.95 15.62 13.22
2048 28.56 15.21 11.43 9.11 7.12 21.23 18.13 15.73 13.27
32  64  134 24.99 13.35 10.35  8.59 7.09 19.61 17.52 15.92 14.21
 128 27.17 14.61 11.27  9.26 7.51 20.99 18.52 16.62 14.59
 256 28.33 15.37 11.83  9.67 7.77 21.80 19.12 17.05 14.81
 512 29.12 15.88 12.20  9.94 7.93 22.33 19.51 17.32 14.94
1024 29.78 16.25 12.47 10.13 8.05 22.71 19.79 17.5 15.03
2048 30.33 16.59 12.72 10.30 8.16 23.07 20.05 17.66 15.11
64  128  269 27.27 14.76 11.47  9.51 7.85 21.25 18.92 17.20 15.40
 256 28.54 15.64 12.15 10.05 8.21 22.24 19.71 17.81 15.76
 512 29.45 16.25 12.62 10.40 8.44 22.88 20.24 18.20 15.97
1024 30.19 16.69 12.96 10.66 8.60 23.35 20.61 18.46 16.10
2048 30.81 17.10 13.27 10.88 8.74 23.79 20.93 18.69 16.21
128  256  538 28.54 15.66 12.19 10.12 8.36 22.28 19.81 18.00 16.16
 512 29.45 16.29 12.69 10.53 8.66 22.96 20.41 18.50 16.48
1024 30.22 16.76 13.07 10.83 8.86 23.47 20.84 18.83 16.68
2048 30.86 17.19 13.41 11.09 9.03 23.95 21.22 19.12 16.84
256  512 1076 29.45 16.29 12.70 10.55 8.71 22.97 20.42 18.54 16.66
1024 30.21 16.76 13.08 10.86 8.95 23.48 20.87 18.92 16.94
2048 30.85 17.20 13.43 11.14 9.15 23.97 21.27 19.26 17.16
512 1024 2152 30.22 16.76 13.08 10.86 8.97 23.48 20.88 18.93 17.00
2048 30.87 17.20 13.43 11.14 9.19 23.97 21.28 19.28 17.28
1024 2048 4303 30.87 17.20 13.43 11.15 9.19 23.97 21.28 19.28 17.29
Funnel Retrieval.
We also designed a simple cascade policy which we call funnel retrieval to successively improve and refine the k-NN shortlist at increasing $D_{s}$. This was an attempt to remove the dependence on manual choice of $D_{s}$ & $D_{r}$. We retrieved a shortlist at $D_{s}$ and then re-ranked the shortlist five times while simultaneously increasing $D_{r}$ (rerank cascade) and decreasing the shortlist length (shortlist cascade), which resembles a funnel structure. We tabulate the performance of funnel retrieval in various configurations in Table 13 on ImageNet-1K, and in Table 14 on ImageNet-4K. With funnel retrieval on ImageNet-1K, we were able to achieve top-1 accuracy within $0.1\%$ of retrieval with $D_{s}$ =2048=2048 (as in Table 8) with a funnel with $D_{s}$ =16=16, with $128\times$ less MFLOPs. Similarly, we are able to achieve equivalent top-1 accuracy within $0.15\%$ of retrieval at $D_{s}$ =2048=2048 (as in Table 10) with funnel retrieval at $D_{s}$ =32=32 on ImageNet-4K, with $64\times$ less MFLOPs. This demonstrates that with funnel retrieval, we can emulate the performance of retrieval with $D_{s}$ =2048=2048 with a fraction of the MFLOPs.
Table 13: Retrieve a shortlist of k-NN with $D_{s}$ sized representations on ImageNet-1K with $\rm MRL$}. This shortlist is then reranked with funnel retrieval, which uses a rerank cascade with a one-to-one mapping with a monotonically decreasing shortlist length as shown in the shortlist cascade. Top-1 and mAP@10 entries (%) within $0.1\%$ of the maximum achievable without reranking on $\rm MRL$} representations, as seen in Table 8, are bolded.
$D_{s}$ Rerank Cascade Shortlist Cascade MFLOPs Top-1 Top-5 Top-10 mAP@10 P@10
[…]
Table 14: Retrieve a shortlist of k-NN with $D_{s}$ sized representations on ImageNet-4K with $\rm MRL$}. This shortlist is then reranked with funnel retrieval, which uses a rerank cascade with a one-to-one mapping with a monotonically decreasing shortlist length as shown in the shortlist cascade. Top-1 and mAP@10 entries (%) within $0.15\%$ of the maximum achievable without reranking on $\rm MRL$} representations, as seen in Table 10, are bolded.
$D_{s}$ Rerank Cascade Shortlist Cascade MFLOPs Top-1 Top-5 Top-10 mAP@10 P@10
[…]
Appendix G Few-shot and Sample Efficiency
We compared $\rm MRL$}, $\rm MRL\text{--}E$}, and FF on various benchmarks to observe the effect of representation size on sample efficiency. We used Nearest Class Means [74] for classification which has been shown to be effective in the few-shot regime [13].
ImageNetV2.
Representations are evaluated on ImageNetV2 with the n-shot k-way setup. ImageNetV2 is a dataset traditionally used to evaluate the robustness of models to natural distribution shifts. For our experiments we evaluate accuracy of the model given $n$ examples from the ImageNetV2 distribution. We benchmark representations in the traditional small-scale (10-way) and large-scale (1000-way) setting. We evaluate for $n\in{1,3,5,7,9}$ with 9 being the maximum value for $n$ because there are 10 images per class.
We observed that $\rm MRL$} had equal performance to FF across all representation sizes and shot numbers. We also found that for both $\rm MRL$} and FF, as the shot number decreased, the required representation size to reach optimal accuracy decreased (Table 15). For example, we observed that 1-shot performance at $32$ representation size had equal accuracy to $2048$ representation size.
Table 15: Few-shot accuracy (%) on ImageNetV2 for 1000-way classification. $\rm MRL$} performs equally to FF across all shots and representation sizes. We also observed that accuracy saturated at a lower dimension for lower shot numbers. E.g. for 1-shot, 32-dim performed comparably to 2048-dim.
Rep. Size Method 1-Shot 3-Shot 5-Shot 7-Shot 9-Shot
8 $F$ 35.41 45.73 49.23 50.89 51.72
$\rm MRL$} 35.37 45.69 49.25 50.85 51.73
16 $F$ 40.88 53.96 57.36 58.72 59.39
$\rm MRL$} 40.90 53.94 57.37 58.65 59.29
32 $F$ 41.41 54.88 58.28 59.63 60.40
$\rm MRL$} 41.40 54.91 58.30 59.65 60.45
64 $F$ 41.25 54.83 58.29 59.82 60.61
$\rm MRL$} 41.28 54.80 58.32 59.77 60.69
128 $F$ 41.36 54.90 58.50 60.05 60.90
$\rm MRL$} 41.38 54.95 58.50 60.06 60.83
256 $F$ 41.36 54.90 58.50 60.05 60.90
$\rm MRL$} 41.38 54.95 58.50 60.06 60.83
512 $F$ 41.36 55.05 58.70 60.19 61.02
$\rm MRL$} 41.34 55.14 58.78 60.40 61.18
1024 $F$ 41.32 55.20 58.85 60.46 61.38
$\rm MRL$} 41.31 55.24 58.86 60.42 61.34
2048 $F$ 41.18 55.09 58.77 60.38 61.34
$\rm MRL$} 41.16 55.10 58.77 60.40 61.28
FLUID.
For the long-tailed setting we evaluated $\rm MRL$} on the FLUID benchmark [87] which contains a mixture of pretrain and new classes. Table 16 shows the evaluation of the learned representation on FLUID. We observed that $\rm MRL$} provided up to 2% higher accuracy on novel classes in the tail of the distribution, without sacrificing accuracy on other classes. Additionally we found the accuracy between low-dimensional and high-dimensional representations was marginal for pretrain classes. For example, the 64-dimensional $\rm MRL$} performed $\sim 1\%$ lower in accuracy compared to the 2048-dimensional counterpart on pretrain-head classes (84.46% vs 85.60%). However for novel-tail classes the gap was far larger (6.22% vs 12.88%). We hypothesize that the higher-dimensional representations are required to differentiate the classes when few training examples of each are known. These results provide further evidence that different tasks require varying capacity based on their difficulty.
Table 16: Accuracy (%) categories indicates whether classes were present during ImageNet pretraining and head/tail indicates classes that have greater/less than 50 examples in the streaming test set. We observed that $\rm MRL$} performed better than the baseline on novel tail classes by $\sim 2\%$ on average.
Rep. Size Method
Pretrain
- Head (>50)
Novel
- Head (>50)
Pretrain
- Tail (<50)
Novel
- Tail (<50)
Mean Per Class
Acc.
Acc.
8 $F$ 68.04 11.30 33.18 0.36 16.29 28.47
$\rm MRL$} 71.75 10.70 38.29 0.19 17.15 29.34
$\rm MRL\text{--}E$} 57.40 6.25 23.14 0.04 11.78 22.81
16 $F$ 80.74 19.12 63.29 2.78 25.65 37.61
$\rm MRL$} 81.79 17.90 61.39 1.95 24.73 37.59
$\rm MRL\text{--}E$} 79.08 9.15 60.33 0.08 20.45 30.24
32 $F$ 83.67 24.30 66.66 4.23 28.86 42.40
$\rm MRL$} 83.46 23.26 65.82 3.75 28.16 41.90
$\rm MRL\text{--}E$} 81.42 10.47 68.01 0.23 22.31 32.17
64 $F$ 84.12 27.49 68.20 5.17 30.64 45.18
$\rm MRL$} 84.46 27.61 67.59 6.22 31.03 45.35
$\rm MRL\text{--}E$} 82.57 13.23 70.18 0.52 23.83 34.74
128 $F$ 84.87 29.96 68.79 5.54 31.84 47.06
$\rm MRL$} 84.88 30.86 68.58 8.41 33.23 47.79
$\rm MRL\text{--}E$} 82.76 18.93 64.46 2.22 25.75 39.19
256 $F$ 84.77 32.78 69.96 7.21 33.65 49.15
$\rm MRL$} 85.10 32.91 69.39 9.99 34.74 49.39
$\rm MRL\text{--}E$} 82.96 22.63 64.55 3.59 27.64 41.96
512 $F$ 85.62 35.27 70.27 9.05 35.42 51.14
$\rm MRL$} 85.62 34.67 70.24 11.43 36.11 50.79
$\rm MRL\text{--}E$} 82.86 25.62 64.34 4.99 29.22 44.20
1024 $F$ 86.30 37.49 71.12 10.92 37.14 52.88
$\rm MRL$} 85.64 35.88 70.02 12.19 36.80 51.58
$\rm MRL\text{--}E$} 83.03 27.78 64.58 6.32 30.57 45.71
2048 $F$ 86.40 37.09 71.74 10.77 37.04 52.67
$\rm MRL$} 85.60 36.83 70.34 12.88 37.46 52.18
$\rm MRL\text{--}E$} 83.01 29.99 65.37 7.60 31.97 47.16
Appendix H Robustness Experiments
Table 17: Top-1 classification accuracy (%) on out-of-domain datasets (ImageNet-V2/R/A/Sketch) to examine robustness of $\rm Matryoshka~Representation~Learning$}. Note that these results are without any fine tuning on these datasets.
ImageNet-V1 ImageNet-V2 ImageNet-R ImageNet-A ImageNet-Sketch
Rep. Size $F$ $\rm MRL\text{--}E$} $\rm MRL$} $F$ $\rm MRL\text{--}E$} $\rm MRL$} $F$ $\rm MRL\text{--}E$} $\rm MRL$} $F$ $\rm MRL\text{--}E$} $\rm MRL$} $F$ $\rm MRL\text{--}E$} $\rm MRL$}
8 65.86 56.92 67.46 54.05 47.40 55.59 24.60 22.98 23.57 2.92 3.63 3.39 17.73 15.07 17.98
16 73.10 72.38 73.80 60.52 60.48 61.71 28.51 28.45 28.85 3.00 3.55 3.59 21.70 20.38 21.77
32 74.68 74.80 75.26 62.24 62.23 63.05 31.28 30.79 31.47 2.60 3.65 3.57 22.03 21.87 22.48
64 75.45 75.48 76.17 63.51 63.15 63.99 32.96 32.13 33.39 2.87 3.99 3.76 22.13 22.56 23.43
128 75.47 76.05 76.46 63.67 63.52 64.69 33.93 33.48 34.54 2.81 3.71 3.73 22.73 22.73 23.70
256 75.78 76.31 76.66 64.13 63.80 64.71 34.80 33.91 34.85 2.77 3.65 3.60 22.63 22.88 23.59
512 76.30 76.48 76.82 64.11 64.09 64.78 35.53 34.20 34.97 2.37 3.57 3.59 23.41 22.89 23.67
1024 76.74 76.60 76.93 64.43 64.20 64.95 36.06 34.22 34.99 2.53 3.56 3.68 23.44 22.98 23.72
2048 77.10 76.65 76.95 64.69 64.17 64.93 37.10 34.29 35.07 2.93 3.49 3.59 24.05 23.01 23.70
Table 18: Zero-shot top-1 image classification accuracy (%) of a ALIGN-$\rm MRL$} model on ImageNet-V1/V2/R/A and ObjectNet.
Rep. Size V1 V2 A R ObjectNet
12 30.57 23.98 14.59 24.24 25.52
24 45.64 37.71 22.75 46.40 35.89
48 53.84 46.16 28.88 60.71 42.76
96 58.31 51.34 33.21 70.12 45.20
192 60.95 53.56 36.10 74.41 48.24
384 62.06 54.77 37.95 76.51 49.10
768 62.26 55.15 37.84 76.73 49.26
Baseline 66.39 59.57 39.97 80.49 51.60
[…]
Appendix I In Practice Costs
All approximate NN search experiments via HNSW32 were run on an Intel Xeon 2.20GHz CPU with 24 cores. All exact search experiments were run with CUDA 11.0 on 2xA100-SXM4 NVIDIA GPUs with 40G RAM each.
$\rm MRL$} models.
As $\rm MRL$} makes minimal modifications to the ResNet50 model in the final fc layer via multiple heads for representations at various scales, it has only an 8MB storage overhead when compared to a standard ResNet50 model. $\rm MRL\text{--}E$} has no storage overhead as it has a shared head for logits at the final fc layer.
Retrieval
Exact search has a search time complexity of $O(dkN)$, and HNSW has a search time complexity of $O(dk\log(N))$, where $N$ is the database size, $d$ is the representation size, and $k$ is the shortlist length. To examine real-world performance, we tabulated wall clock search time for every query in the ImageNet-1K and ImageNet-4K validation sets over all representation sizes $d$ in Table 19 for both Exact Search and HNSW32, and ablated wall clock query time over shortlist length $k$ on the ImageNet-1K validation set in Table 21. The wall clock time to build the index and the index size is also shown in Table 20.
Table 19: Retrieval k-NN wall clock search times (s) over the entire validation (query) set of ImageNet-1K and ImageNet-4K, containing 50K and 200K samples respectively.
Rep. Size ImageNet-1K ImageNet-4K
ExactL2 HNSW32 ExactL2 HNSW32
8 0.60 0.14  35.70  1.17
16 0.57 0.18  36.16  1.65
32 0.60 0.20  36.77  1.75
64 0.66 0.24  27.88  2.21
128 0.86 0.32  30.10  4.15
256 1.29 0.46  34.97  3.39
512 2.17 0.68  46.97  4.83
1024 3.89 1.05  70.59  7.14
2048 7.31 2.05 117.78 13.43
Table 20: FAISS [45] index size and build times for exact k-NN search with L2 Distance metric and approximate k-NN search with HNSW32 [59].
Rep. Size Exact Search HNSW32
ImageNet-1K ImageNet-4K ImageNet-1K ImageNet-4K
Index Size
Index Build
Time (s)
Index Size
Index Build
Time (s)
Index Size
Index Build
Time (s)
Index Size
Index Build
Time (s)
8    40  0.04   131  0.33   381  4.87  1248  24.04
16    80  0.08   263  0.27   421  6.15  1379  33.31
32   160  0.16   525  0.52   501  6.80  1642  37.41
64   320  0.38  1051  1.05   661  8.31  2167  47.23
128   641  0.64  2101  2.10   981 11.73  3218  89.87
256  1281  1.27  4202  4.20  1622 17.70  5319 102.84
512  2562  2.52  8404  8.39  2903 27.95  9521 158.47
1024  5125  5.10 16808 17.20  5465 44.02 17925 236.30
2048 10249 10.36 33616 41.05 10590 86.15 34733 468.18
Table 21: Retrieval k-NN wall clock search times (s) over entire validation (query) set of ImageNet-1K over various shortlist lengths $k$.
Index k = 50 k = 100 k = 200 k = 500 k = 1000 k = 2048
Exact L2 0.4406 0.4605 0.5736 0.6060 1.2781 2.7047
HNSW32 0.1193 0.1455 0.1833 0.2145 0.2333 0.2670
Appendix J Analysis of Model Disagreement
Class Trends
Does increasing representation size necessarily help improve classification performance across all classes in ImageNet-1K? We studied this question by examining trends in performance with increasing representation size from d=8,2048d={8,...2048}. For $\rm MRL$} models, we observed that $244$ classes showed a monotonic improvement in performance with increasing $d$, $177$ classes first improved but then observed a slight dip (one or two misclassifications per class), $49$ classes showed a decline first and then an improvement, and the remaining classes did not show a clear trend. When we repeated this experiment with independently trained FF models, we noticed that $950$ classes did not show a clear trend. This motivated us to leverage the disagreement as well as gradual improvement of accuracy at different representation sizes by training $\rm Matryoshka~Representations$}. Figure 12 showcases the progression of relative per-class accuracy distribution compared to the $\rm Matryoshka~Representation~Learning$}-2048 dimensional model. This also showed that some instances and classes could benefit from lower-dimensional representations.
Discussion of Oracle Accuracy
Based on our observed model disagreements for different representation sizes $d$, we defined an optimal oracle accuracy [56] for $\rm MRL$}. We labeled an image as correctly predicted if classification using any representation size was correct. The percentage of total samples of ImageNet-1K that were firstly correctly predicted using each representation size $d$ is shown in Table 22. This defined an upper bound on the performance of $\rm MRL$} models, as $18.46\%$ of the ImageNet-1K validation set were incorrectly predicted $\forall d\in\{8,16,\ldots,2048\}$. We show the oracle performance on $\rm MRL$} models for ImageNet-1K/V2/A/R/Sketch datasets in Table 23.
Figure 12: Progression of relative per-class accuracy vs $\rm MRL$}-2048. As the dimensionality increases, the spread shrinks while the class marked (x) (Madagascar cat) loses accuracy.
In an attempt to derive an optimal routing policy to emulate oracle accuracy, we designed the adaptive classification via cascading method as discussed in Appendix D.1. This led to an interesting observation on the expected dimensionality for $76.30\%$ top-1 classification accuracy being just $d\sim 37$. We leave the design and learning of a more optimal policy for future work.
Table 22: Percentage of ImageNet-1K validation set that is first correctly predicted using each representation size $d$. We note that $18.46\%$ of the samples cannot be correctly predicted by any representation size. The remaining $81.54\%$ constitutes the oracle accuracy.
Rep. Size 8 16 32 64 128 256 512 1024 2048
Always
Wrong
Correctly
Predicted
67.46 8.78 2.58 1.35 0.64 0.31 0.20 0.12 0.06 18.46
[…]
Top-1 ImageNetV1 ImageNetV2 ImageNet-A ImageNet-R ImageNet-Sketch
FF–2048 76.9 64.9 3.6 35.1 23.7
$\rm MRL$}–Oracle 81.5 70.6 8.7 39.8 28.9
Grad-CAM Examples
We analyzed the nature of model disagreement across representation sizes with $\rm MRL$} models with the help of Grad-CAM visualization [75]. We observed there were certain classes in ImageNet-1K such as "tools", "vegetables" and "meat cutting knife" which were occasionally located around multiple objects and a cluttered environment. In such scenarios, we observed that smaller representation size models would often get confused due to other objects and fail to extract the object of interest which generated the correct label. We also observed a different nature of disagreement arising when the models got confused within the same superclass. For example, ImageNet-1K has multiple "snake" classes, and models often confuse a snake image for an incorrect species of snake.
Superclass Performance
We created a 30 superclass subset of the validation set based on wordnet hierarchy (Table 24) to quantify the performance of $\rm MRL$} model on ImageNet-1K superclasses. Table 25 quantifies the performance with different representation size.
Table 24: 30 Superclasses in ImageNet-1K corresponding to the performance in Table 25.
insect motor vehicle artiodactyl vegetable game equipment
terrier serpent machine measuring device sheepdog
protective covering sporting dog vessel, watercraft building lizard
garment hound monkey home appliance wind instrument
vessel fish nourishment electronic equipment oscine
furniture wading bird tool canine mechanism
Table 25: Performance of $\rm MRL$} model on 31-way classification (1 extra class is for reject token) on ImageNet-1K superclasses.
Rep. Size 8 16 32 64 128 256 512 1024 2048
$\rm MRL$} 85.57 88.67 89.48 89.82 89.97 90.11 90.18 90.22 90.21
Appendix K Ablation Studies
K.1 $\rm MRL$} Training Paradigm
Table 26: Top-1 classification accuracy (%) on ImageNet-1K of various ResNet50 models which are finetuned on pretrained FF-2048 model. We observed that adding more non-linearities is able to induce nesting to a reasonable extent even if the model was not pretrained with nesting in mind.
Rep. Size fc
4.2 conv3,
fc
4.2 conv2,
conv3, fc
4.2 full,
fc
All ($\rm MRL$})
8  5.15 36.11 54.78 60.02 66.63
16 13.79 58.42 67.26 70.10 73.53
32 32.52 67.81 71.62 72.84 75.03
64 52.66 72.42 73.61 74.29 75.82
128 64.60 74.41 74.67 75.03 76.30
256 69.29 75.30 75.23 75.38 76.47
512 70.51 75.96 75.47 75.64 76.65
1024 70.19 76.18 75.70 75.75 76.76
2048 69.72 76.44 75.96 75.97 76.80
$\rm Matryoshka~Representations$} via Finetuning.
To observe if nesting can be induced in models that were not explicitly trained with nesting from scratch, we loaded a pretrained FF-2048 ResNet50 model and initialized a new $\rm MRL$}  layer, as defined in Algorithm 2, Appendix C. We then unfroze different layers of the backbone to observe how much non-linearity in the form of unfrozen conv layers needed to be present to enforce nesting into a pretrained FF model. A description of these layers can be found in the ResNet50 architecture [27]. All models were finetuned with the FFCV pipeline, with same training configuration as in the end-to-end training aside from changing lr =0.1=0.1 and epochs =10=10. We observed that finetuning the linear layer alone was insufficient to learn $\rm Matryoshka~Representations$} at lower dimensionalities. Adding more and more non-linear conv+ReLU layers steadily improved classification accuracy of $d=8$ from $5\%$ to $60\%$ after finetuning, which was only $6\%$ less than training $\rm MRL$} end-to-end for 40 epochs. This difference was successively less pronounced as we increased dimensionality past $d=64$, to within $1.5\%$ for all larger dimensionalities. The full results of this ablation can be seen in Table 26.
Table 27: An ablation over boosting training loss at lower nesting dimensions, with top-1 and top-5 accuracy (%). The models are described in Appendix K.1.
Model $\rm MRL$} $\rm MRL$}-8boost $\rm MRL$}-8+16boost
Rep. Size Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
8 66.63 84.66 69.53 86.19 69.24 85.96
16 73.53 89.52 73.86 89.44 73.91 89.55
32 75.03 91.31 75.28 91.21 75.10 91.14
64 75.82 92.27 75.84 92.22 75.67 92.06
128 76.30 92.82 76.28 92.74 76.07 92.52
256 76.47 93.02 76.48 92.97 76.22 92.72
512 76.65 93.13 76.56 93.09 76.35 92.85
1024 76.76 93.22 76.71 93.21 76.39 92.98
2048 76.80 93.32 76.76 93.28 76.52 93.05
Relative Importance.
We performed an ablation of $\rm MRL$} over the relative importance, $c_{m}$, of different nesting dimensions $m\in\cal{M}$, as defined in Sec. 3. In an attempt to improve performance at lower dimensionalities, we boosted the relative importance $c_{m}$ of training loss at lower dimensions as in Eq. 1 with two models, $\rm MRL$}-8boost and $\rm MRL$}-8+16boost. The $\rm MRL$}-8boost model had $c_{m\in\cal M}=[2,1,1,1,1,1,1,1,1]$ and the $\rm MRL$}-8+16boost model had $c_{m\in\cal M}=[2,1.5,1,1,1,1,1,1,1]$. The relative importance list $c_{m\in\cal M}$ had a 1-to-1 correspondence with nesting dimension set $\mathcal{M}$. In Table 27, we observed that $\rm MRL$}-8boost improves top-1 accuracy by $3\%$ at $d=8$, and also improves top-1 accuracy of all representation scales from 16 to 256 over $\rm MRL$}, while only hurting the performance at 512 to 2048 representation scales by a maximum of 0.1%. This suggests that the relative importance $c_{m}$ can be tuned/set for optimal accuracy for all $m\in\mathcal{M}$, but we leave this extension for future work.
$\rm Matryoshka~Representations$} at Arbitrary Granularities.
To train $\rm MRL$}, we used nested dimensions at logarithmic granularities $\mathcal{M}=\{8,16,\ldots,1024,2048\}$ as detailed in Section 3. We made this choice for two empirically-driven reasons: a) The accuracy improvement with increasing representation size was more logarithmic than linear (as shown by FF models in Figure 3). This indicated that optimizing for granularities increasing in a non-logarithmic fashion would be sub-optimal both for maximum performance and expected efficiency; b) If we have $m$ arbitrary granularities, the expected cost of the linear classifier to train $\rm MRL$} scales as $O(L*(m^{2}))$ while logarithmic granularities result in O(L2log(d))O(L*2log(d)) space and compute costs.
To demonstrate this effect, we learned $\rm Matryoshka~Representations$} with uniform ($\rm MRL$}-Uniform) nesting dimensions $m\in\mathcal{M}=\{8,212,416,620,824,1028,1232,1436,1640,1844,2048\}$. We evaluated this model at the standard ($\rm MRL$}-log) dimensions $m\in\mathcal{M}=\{8,16,32,64,128,256,512,1024,2048\}$ for ease of comparison to reported numbers using 1-NN accuracy (%). As shown in Table 29, we observed that while performance interpolated, $\rm MRL$}-Uniform suffered at low dimensions as the logarithmic spacing of $\rm MRL$}-log resulted in tighter packing of information in these initial dimensions. The higher nesting dimensions of $\rm MRL$}-Uniform did not help in significant accuracy improvement due to accuracy saturation, which is often logarithmic in representation size as shown by FF models. Note that the slight improvement at dimensions higher than 512 for $\rm MRL$}-Uniform is due to multiple granularities around them compared to just three for $\rm MRL$}-log, which are not useful in practice for efficiency.
Lower Dimensionality.
We experimented with training $\rm MRL$} with smaller nesting dimension than $m=8$, as shown in Table 29, with two models: MRL-4 and MRL-6. We found that using lower than 8-dimensions to train $\rm MRL$}, i.e. $m_{0}\in\{4,6\}$ for MRL-4 and MRL-6 respectively, did not affect the top-1 accuracy of other granularities significantly. However, granularities smaller than 8-dimensions had very low accuracy and were often unusable for deployment along with additional training difficulty. We also observed a small dip in accuracy at higher dimensions which we attribute to the joint loss that now also included the harder optimization of the smallest dimension. Lastly, we hypothesize the dimensionality of 8 is an empirically validated design choice due to the considerable accuracy it provided along with the ease of training.
Table 28: An ablation over training with smaller nesting dimensionalities in terms of Top-1 accuracy (%). MRL-4 and MRL-6 are variations of the original model (MRL-8) with $m_{0}\in\{4,6\}$, where $m\in\mathcal{M}$ is part of the nesting_list as seen in Alg 2.
Rep. Size MRL-4 MRL-6 MRL-8
4 27.25 - -
6 - 58.71 -
8 66.86 67.55 66.63
16 73.36 73.10 73.53
32 74.82 74.49 75.03
64 75.51 75.32 75.82
128 75.93 75.61 76.30
256 76.08 75.82 76.47
512 76.31 75.93 76.65
1024 76.38 76.04 76.76
2048 76.43 76.12 76.80
Table 29: An ablation over training $\rm MRL$} with nesting list at uniformly distributed granularities. Entries in the $\rm MRL$}-Uniform column are evaluated at logarithmic dimensions for a fair comparison to $\rm MRL$}-Log (standard $\rm MRL$}) with 1-NN accuracy (%).
Rep. Size $\rm MRL$}-Log $\rm MRL$}-Uniform
8 62.19 58.44
16 67.91 61.11
32 69.46 63.82
64 70.17 66.44
128 70.52 68.71
256 70.62 70.06
512 70.82 70.98
1024 70.89 71.37
2048 70.97 71.44
K.2 Retrieval
Adaptive Retrieval.
To examine the effect of increasing shortlist lengths on search time, we performed a reranking ablation over shortlist lengths for $D_{s}$= 16 and $D_{r}$= 2048 over ImageNet-1K  in Table 30, and over ImageNet-4K in Table 31. We observed that using a larger shortlist $k$ saturated ImageNet-1K performance at kk=200. But using larger shortlists until $k=2048$, the maximum value supported by the FAISS framework, steadily improved performance on ImageNet-4K. This is likely due to the increased database size, but could also indicate a correlation with ImageNet-4K being slightly out-of-distribution making the task at hand harder.
Table 30: Adaptive retrieval ablation over shortlist length $k$ for $D_{s}=16$, $D_{r}=2048$ on ImageNet-1K  with exact search. Entries with the highest P@1 and mAP@10 across all $k$ are in bold.
Shortlist
Length
P@1 mAP@10 mAP@25 mAP@50 mAP@100 P@10 P@25 P@50 P@100
 100 70.88 65.19 63.62 62.59 61.24 69.96 69.24 68.53 67.20
 200 70.90 65.27 63.73 62.82 61.97 70.10 69.44 68.90 68.21
 400 70.94 65.26 63.71 62.81 62.03 70.15 69.51 69.02 68.47
 800 70.96 65.23 63.64 62.69 61.85 70.16 69.52 69.02 68.45
1600 70.96 65.20 63.58 62.58 61.66 70.16 69.5 68.97 68.36
2048 70.97 65.20 63.57 62.58 61.64 70.16 69.5 68.97 68.35
Table 31: Adaptive retrieval ablation over shortlist length $k$ for $D_{s}=16$, $D_{r}=2048$ on ImageNet-4K with exact search.
Shortlist
Length
P@1 mAP@10 mAP@25 mAP@50 mAP@100 P@10 P@25 P@50 P@100
 100 27.70 14.38 10.62  8.26 6.07 20.12 16.87 14.29 11.26
 200 28.56 15.21 11.43  9.11 7.12 21.23 18.13 15.73 13.27
 400 29.34 15.83 12.06  9.76 7.79 22.08 19.09 16.83 14.54
 800 29.86 16.30 12.53 10.23 8.26 22.72 19.83 17.65 15.45
1600 30.24 16.63 12.86 10.56 8.60 23.18 20.36 18.23 16.11
2048 30.35 16.73 12.96 10.65 8.69 23.31 20.50 18.40 16.30